### 31 Mar 2026 v2

# Closed-form finite-time blow-up and conditional perturbative stability framework for a (1+2)(1+2)D system (E1) derived from the 2D inviscid Boussinesq equations

Abstract.
In polar variables on a planar wedge, we study a closed (1+2)(1+2)D system (E1) derived from the two-dimensional inviscid Boussinesq equations under a parity ansatz. The system retains a stretching-type interaction between scalar building blocks while remaining explicit enough to support both closed-form blow-up profiles and a weighted remainder analysis.
The paper has three main outputs. First, we identify the special ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting explicit finite-time blow-up mechanism. Second, after embedding these ridge profiles into the full wedge by anisotropic seeds, we derive the exact remainder equations around a prescribed background and prove singular weighted linear estimates for the remainder system. Third, we formulate a conditional nonlinear control principle in the spirit of Elgindi–Jeong: if the background exists on $[0,T)$ with the coefficient bounds required by the weighted energy method, and if the remainder remains subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
What is unconditional in the present revision is the exact reduction from Boussinesq, the explicit ridge blow-up mechanism, and the weighted remainder framework. What remains conditional is the full-wedge background construction up to the blow-up time together with the rigidity properties needed to close the bootstrap without loss. In this sense, the manuscript isolates the background existence problem as the main remaining step toward a complete nonlinear stability theorem for the blow-up scenario.
Key words and phrases:
2D inviscid Boussinesq, Constantin–Lax–Majda type model, finite-time blow-up, explicit solutions, conditional perturbative framework, ridge reduction, weighted Sobolev energy, divergence-form closure
2020 Mathematics Subject Classification:
35B44, 35B40, 35Q86, 76B03, 76D05
2020 Mathematics Subject Classification. Primary 35B44, 35Q86, 76B03; Secondary 35B35, 35Q35.
Keywords. 2D inviscid Boussinesq equations; Constantin–Lax–Majda type model; explicit blow-up solutions; ridge reduction; weighted energy; linear estimates; conditional nonlinear control; weighted Sobolev energies.
1. Introduction
The question of whether smooth solutions can develop a singularity in finite time is a central theme in nonlinear PDE and fluid mechanics. In this paper we study a $(1+2)$-dimensional closed subsystem (E1), rigorously derived from the inviscid 2D Boussinesq equations: it retains a stretching-like interaction between two scalar fields while suppressing enough geometric and nonlocal complexity to permit closed-form analysis. Our aim is twofold: to construct explicit finite-time blow-up backgrounds and to formulate a perturbative stability theory around those backgrounds that is mathematically solid at the linear level and explicit about the remaining nonlinear obstruction.
We will call the Hou-Li type new variables $\{u,v,g\}$ of  (2.2) the building blocks of vorticity, because their units are equal to the unit of $\boldsymbol{\omega}=\nabla\times\boldsymbol{u}$. Also the quadratic vortex stretching terms are greatly simplified: $(uv,\,v^{2}-u^{2},\,-g^{2})$.
Related work and context.
As emphasized by Elgindi–Jeong [17] and Elgindi–Pasqualotto [19], singularity formation for the inviscid 2D Boussinesq and 3D Euler systems has a long history. We recall only a small selection of representative references here, emphasizing works most closely aligned with the closed-form mechanism and stability framework developed below. For the inviscid 2D Boussinesq system, classical local theory and conditional breakdown criteria go back to Cannon–DiBenedetto [12], Chae–Nam [3] (see also Chae–Kim–Nam [7]), and Taniuchi [29]; see also Wu’s lecture notes [30] and the Euler perspective of Beale–Kato–Majda [1] and Constantin [10, 11]. On the modeling side, explicit/didactic mechanisms include Chae–Constantin–Wu [2] and the CLM/De Gregorio lineage [9, 16], as well as 1D reductions for Boussinesq-type dynamics such as Choi–Kiselev–Yao [8]. For rigorous singularity constructions and perturbative stability scenarios in PDE settings, see Elgindi–Jeong [18, 17], Chen–Hou [4, 5], and the synthesis of Drivas–Elgindi [15].
Main achievements.
  • (0)
    We derive the closed subsystem (E1) exactly from the inviscid 2D Boussinesq equations under a parity ansatz and identify the variables $\{u,v,g\}$ as convenient vorticity building blocks.
  • (1)
    We identify the ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type and record the resulting explicit finite-time blow-up profiles.
  • (2)
    We derive the exact remainder equations around a prescribed background in the updated $(x,\xi)$ variables, with all pure-background contributions retained in the background system.
  • (3)
    We prove weighted singular linear estimates and formulate a conditional nonlinear remainder theorem of Elgindi type: once the background is available with the required coefficient bounds, blow-up transfers from the background to the full solution.
  • (4)
    We isolate the remaining open step in the program, namely the full-wedge background construction up to the blow-up time together with the compatibility structure needed to close the nonlinear bootstrap.
Organization. Section 2 derives the closed subsystem (2.5) from the inviscid 2D Boussinesq equations and identifies the ridge rays on which (E1) reduces to a (1+1)(1+1)D CLM-type reaction system. Section 2.6 rewrites the background and remainder system in the updated $(x,\xi)$ variables and records the exact remainder equations. Section 4 collects the explicit background formulas and the corresponding weighted pointwise and energy bounds. Section 5 proves the singular weighted linear estimates, formulates the conditional nonlinear remainder mechanism, and explains how blow-up transfers from the background to the full solution once the remaining background hypotheses are verified.
2. The derivation of system (E1) from 2D inviscid Boussinesq equations
2.1. The derivation of system (E1) from 2D inviscid Boussinesq equations
In this section we rewrite the velocity–pressure form of the 2D inviscid Boussinesq equations (see, for example, Wu [30], Elgindi–Jeong [18], and Kiselev–Pu–Yao [24]) in terms of new variables that we call vorticity building blocks. In these variables, the stretching and transport structure becomes more transparent.
In the velocity-pressure form, the 2D inviscid Boussinesq equations for velocity $\boldsymbol{u}=u_{2}\boldsymbol{e}_{2}+u_{3}\boldsymbol{e}_{3}$, pressure $P$ and density $\vartheta$ (assuming the gravity in $\boldsymbol{e}_{2}$ direction) in $\mathbb{R}^{2}$ are given by
[…]
Assume that $u_{2}$ is odd in $x_{2}$ and even in $x_{3}$, that $u_{3}$ is even in $x_{2}$ and odd in $x_{3}$, that $p$ is even in $(x_{2},x_{3})$, and that $\vartheta$ is odd in $x_{2}$ and even in $x_{3}$. Define the Hou–Li [23] type variables
$\{v,g,u^{2},p\}:=\left\{-\frac{u_{2}}{x_{2}},\frac{u_{3}}{x_{3}},\frac{\vartheta}{x_{2}},P\right\}$ (2.2)
Then  (2.1) can be converted into the following system (with $(\rho,z)=(x_{2},x_{3})$ for notational convenience):
Verification of the reduction. Under the stated parity assumptions, the whole plane system (2.1) can be extended to $\mathbb{R}^{2}$ by odd/even reflection. In particular, along the symmetry axes $x_{2}=0$ and $x_{3}=0$ the quotients
$v:=-\frac{u_{2}}{x_{2}},\qquad g:=\frac{u_{3}}{x_{3}},\qquad u^{2}:=\frac{\vartheta}{x_{2}}$ (2.3)
extend smoothly (the numerator has the matching odd symmetry), and we may write
$u_{2}=-\rho\,v,\qquad u_{3}=z\,g,\qquad\vartheta=\rho\,u^{2}.$
With this convention the material derivative becomes
$\tfrac{\mathrm{D}}{\mathrm{D}t}=\partial_{t}+u_{2}\partial_{\rho}+u_{3}\partial_{z}=\partial_{t}-\rho v\,\partial_{\rho}+zg\,\partial_{z},$ (2.4)
which matches the last line in (2.5). Now:
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}\vartheta=0$ and $\vartheta=\rho u^{2}$,
    $0=\tfrac{\mathrm{D}}{\mathrm{D}t}(\rho u^{2})=\Big(\tfrac{\mathrm{D}\rho}{\mathrm{D}t}\Big)u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=u_{2}u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=-\rho v\,u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2}),$
    hence $\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=vu^{2}$, equivalently $\tfrac{\mathrm{D}}{\mathrm{D}t}u=\frac{1}{2}uv$.
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}u_{2}=-\partial_{\rho}P+\vartheta$ and $u_{2}=-\rho v$,
    $\tfrac{\mathrm{D}}{\mathrm{D}t}u_{2}=\tfrac{\mathrm{D}}{\mathrm{D}t}(-\rho v)=-(\tfrac{\mathrm{D}\rho}{\mathrm{D}t})v-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t}=-u_{2}v-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t}=\rho v^{2}-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t},$
    so dividing by $\rho>0$ gives $\tfrac{\mathrm{D}v}{\mathrm{D}t}=v^{2}-u^{2}+\frac{1}{\rho}P_{\rho}$.
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}u_{3}=-\partial_{z}P$ and $u_{3}=zg$,
    $\tfrac{\mathrm{D}}{\mathrm{D}t}u_{3}=\tfrac{\mathrm{D}}{\mathrm{D}t}(zg)=(\tfrac{\mathrm{D}z}{\mathrm{D}t})g+z\tfrac{\mathrm{D}g}{\mathrm{D}t}=u_{3}g+z\tfrac{\mathrm{D}g}{\mathrm{D}t}=zg^{2}+z\tfrac{\mathrm{D}g}{\mathrm{D}t},$
    hence $\frac{\mathrm{D}g}{\mathrm{D}t}=-g^{2}-\frac{1}{z}P_{z}$.
  • Finally, incompressibility $\partial_{\rho}u_{2}+\partial_{z}u_{3}=0$ with $u_{2}=-\rho v$, $u_{3}=zg$ yields
    $0=\partial_{\rho}(-\rho v)+\partial_{z}(zg)=-(v+\rho v_{\rho})+(g+zg_{z}),$
    i.e. $zg_{z}-\rho v_{\rho}+g-v=0$.
This proves that the change of variables reduces (2.1) exactly to (2.5).
[…]
Remark 2.1.
Inspection of  (2.5) shows that if the initial data for $\{u,v,g,p\}$ are symmetric in both $r$ and $z$, then the PDE system preserves these symmetries. In this sense,  (2.5) may be regarded as posed on $\mathbb{R}^{2}$.
Remark 2.2.
We will call $\{u,v,g\}$ the building blocks of vorticity, because their units are equal to the unit of $\boldsymbol{\omega}=\nabla\times\boldsymbol{u}$. Also the quadratic vortex stretching terms are greatly simplified: $(uv,\ v^{2}-u^{2},\ -g^{2})$.
2.2. System (E11)
We now introduce a stream function $\bar{\psi}$ and augment  (2.5) with two additional relations, obtaining system  (2.6). To reserve the symbols $(u,v,g,p)$ for later perturbation variables, we place bars on the background unknowns. Thus system (E11) in  (2.6) consists of five dependent variables $(\bar{u},\bar{v},\bar{g},\bar{p},\bar{\psi})$, viewed as even functions of $(\rho,z)$ on the meridian plane, together with six equations; the last equation defines $\tfrac{\mathrm{D}}{\mathrm{D}t}$.
$\text{(i) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{u}-\tfrac{1}{2}\,\bar{u}\,\bar{v},\quad(t,\rho,z)\in[0,T)\times\mathbb{R}^{2}$ (2.6a)
$\text{(ii) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{v}-\bar{v}^{2}+\bar{u}^{2}-\tfrac{1}{\rho}\bar{p}_{\rho},$ (2.6b)
$\text{(iii) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{g}+\bar{g}^{2}+\tfrac{{\mu^{2}}}{z}\,\,\bar{p}_{z},$ (2.6c)
$\text{(iv) }0$ $=z\partial_{z}\bar{g}-\rho\partial_{\rho}\bar{v}+\bar{g}-\bar{v},$ (2.6d)
$\text{(v) }0$ $=\bar{v}-\bar{\psi}-z\partial_{z}\bar{\psi},$ (2.6e)
$\text{(vi) }0$ $=\bar{g}-\bar{\psi}-\rho\partial_{\rho}\bar{\psi},$ (2.6f)
$\text{(vii) }\tfrac{\mathrm{D}}{\mathrm{D}t}:$ $={{\color[rgb]{0,0,1}\lambda}}\partial_{t}+\bar{g}\,z\partial_{z}-\bar{v}\,\rho\partial_{\rho},.$ (2.6g)
[…]
We will study the system in polar coordinates.
2.3. Polar coordinates $(x=r=\sqrt{\rho^{2}+z^{2}},\ \theta=\arctan(z/\rho))$
We use polar coordinates on the meridian plane:
$\rho=x\cos\left(\theta\right),\quad z=x\sin\left(\theta\right).$ (2.7)
Remark 2.4.
The polar coordinates $(x,\theta)$ on the meridian plane $(\rho,z)$ are also the spherical coordinates $(x,\theta,\phi)$ (with north pole at $\theta=\pi/2$) for 3D axisymmetric functions in $\mathbb{R}^{3}$.
2.4. Background
We write the background solutions as
$\bar{u}$ $=U(t,x,\xi),\quad\bar{v}=V(t,x,\xi),\quad\bar{g}=G(t,x,\xi),$ (2.8)
$\bar{p}$ $=P(t,x,\xi),\quad\xi=\tan(\theta).$
After substituting  (3.1) into  (2.6),we obtain four equations with the following structure:
$\left\{\begin{aligned} \lambda U_{t}-\tfrac{1}{2}V\,U&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xU_{x}-(G+V)\xi U_{\xi},\\[5.69054pt] \lambda V_{t}-V^{2}+U^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xV_{x}-(G+V)\xi V_{\xi}+\tfrac{1}{x}P_{x}-\tfrac{1+\xi^{2}}{x^{2}}\xi{P_{\xi}},\\[5.69054pt] \lambda G_{t}+G^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xG_{x}-\tfrac{1}{2}\xi(G^{2}+V^{2})_{\xi}-\tfrac{\mu}{x}P_{x}-\tfrac{\mu(1+\xi^{2})}{x^{2}\xi^{2}}\xi{P_{\xi}},\\[5.69054pt] G-V&=-\tfrac{1}{1+\xi^{2}}(\xi^{2}xG_{x}-xV_{x})+\xi(G_{\xi}+V_{\xi}).\end{aligned}\right$. (2.9)
We now examine how the equations simplify under the following Ansatz (I):
$\xi=\xi_{0}=\pm 1,\quad G(t,x,\xi_{0})=V(t,x,\xi_{0}).$ (2.10)
and ridge-flat Ansatz (II) (ridge-flatness condition in the directions normal to the ridge):
$(V_{\xi},U_{\xi},G_{\xi},P_{\xi})|_{\xi_{0}}=0,$ (2.11)
2.5. Ridge ray and Ridge functions
Theorem 2.5.
System  (2.6)(including the divergence-free condition  (2.6)(4)) restricted to the rays determined by $\xi^{2}=\xi_{0}^{2}=1$
Assuming Ansatz (I) and Ansatz (II), then we have:
(A): Divergence constraint and ridge flatness fixes the ridge rays. Under (2.10) and (2.11), the divergence identity (2.6d) implies
$\xi_{0}=\tan(\theta_{0})=\pm 1.$
Equivalently,
$\theta_{0}\in\Bigl\{\pm\tfrac{\pi}{4},\ \pi\pm\tfrac{\pi}{4}\Bigr\}.$
In the (r,z)(r,z)–plane this corresponds to the two straight lines through the origin
$z=\pm\,r,$
i.e. four rays (two rays on each diagonal line). When working in the first quadrant $r\geq 0$, $z\geq 0$, we take the principal choice
$\theta_{0}=\tfrac{\pi}{4}.$
(B). Convection-free on the principal ridge ray. On $\xi=\xi_{0}$,
$\left.\tfrac{\mathrm{D}}{\mathrm{D}t}\right|_{\xi=\xi_{0}}=\partial_{t}.$
(C) With the specific choice of scaling parameters $(\lambda,\mu)$:
$\lambda=\tfrac{3}{2},\quad\mu=\left(\tfrac{5}{3}\right)^{1/2},$ (2.12)
the dynamics of all ridge functions $\{U,V,G,P\}(t,x,\theta_{0})$ are completely determined by the following $1+1$-dimensional convection-free reaction system for $U(t,x,\xi_{0})$ and $V(t,x,\xi_{0})$.
$\left\{\begin{aligned} &U_{t}=\tfrac{1}{3}\,V\,U,\\[5.69054pt] &V_{t}=\tfrac{1}{6}V^{2}-\tfrac{5}{12}U^{2}.\end{aligned}\right$. (2.13)
Notice that (2.13) is point-wise ODE system. It holds for every point $x\in\mathbb{R}$.
After setting $\tau=6t$, $\left(\tfrac{2}{5}\right)^{1/2}U(t,x)=\tfrac{1}{2}\omega(\tau,x)$, and $V(t,x)=\tfrac{1}{2}H\left(\omega(\tau,x)\right)$, we see that this system becomes algebraically equivalent, at each fixed $x$, to the Constantin–Lax–Majda reaction ODE system proposed as a model for vorticity stretching.
$\left\{\begin{aligned} &\partial_{\tau}\omega=\omega H(\omega)\\[5.69054pt] &\partial_{\tau}H(\omega)=\tfrac{1}{2}\left(H(\omega)^{2}-\omega^{2}\right).\end{aligned}\right$. (2.14)
Proof of theorem 2.5.
Claim (A) follows directly from  (3.2).
Because $\xi_{0}^{2}=1$ and $G(t,x,\xi_{0})=V(t,x,\xi_{0})$ on the ridge rays, the convection terms $(\xi^{2}G-V)\partial_{x}$ vanish in  (3.2). This proves Claim (B).
Thus our ridge ray system is composed of 4 PDEs, as shown below
$\left\{\begin{aligned} \tfrac{3}{2}U_{t}&=\tfrac{1}{2}V\,U,\\[5.69054pt] \tfrac{3}{2}V_{t}&=V^{2}-U^{2}+\tfrac{1}{x}P_{x},\\[5.69054pt] \tfrac{3}{2}V_{t}&=-V^{2}-\tfrac{5}{3x}P_{x}.\end{aligned}\right$. (2.15)
Separating $P_{x}$ from $V_{t}$ yields  (2.13) together with the simplified Poisson equation for $P(t,x,\xi_{0})$.
$P_{x}$ $=\tfrac{3x}{8}\left(U^{2}-2V^{2}\right),$ (2.16)
This proves Claim (C) and completes the Proof of theorem 2.5. ∎
It is well known that system  (2.13), like the Constantin–Lax–Majda system, admits the following closed-form solutions.
[…]
If we prescribe the following initial conditions
$\left\{\begin{aligned} &r=x^{2}+A_{1}(\xi^{2}-1)^{2m},\\ &a=a(x)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b=b(x)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0.\end{aligned}\right$. (2.18)
Then this solution blows up at $(t,\xi)=(0,\pm 1)$ in finite time $T=\tfrac{6}{A}$.
$V(t,x,\pm 1)\to\infty,\quad\tfrac{U(t,x,\pm 1)}{U(0,x,\pm 1)}\to\infty,\text{ as }t\to T=\tfrac{6}{A},x\to 0.$ (2.19)
2.6. Ridge blow-up and the full-wedge extension problem
The explicit formulas (2.17) immediately produce a blowing-up ridge background. What remains open is the extension of that ridge profile to a full classical solution of (2.9) on the wedge $x>0$, $|\xi|\leq 1$ while preserving enough ridge flatness to keep the reduced dynamics exact.
Theorem 2.6 (Explicit ridge background and localized finite-time blow-up).
Let
$a_{\mathrm{ridge}}(x):=\frac{A}{(1+x^{2})^{3}},\qquad b_{\mathrm{ridge}}(x):=\frac{B\,x^{2}}{(1+x^{2})^{6}},\qquad A,B>0,$
and let $(V_{\mathrm{ridge}},U_{\mathrm{ridge}})$ be defined by (2.17) with $(a,b)=(a_{\mathrm{ridge}}(x),b_{\mathrm{ridge}}(x))$. Then:
  1. (1)
    for every fixed $x\in\mathbb{R}$, the pair $(V_{\mathrm{ridge}}(t,x),U_{\mathrm{ridge}}(t,x))$ is smooth on $[0,T)$, where
    $T=\frac{6}{A};$
  2. (2)
    for every fixed $x\neq 0$, both $V_{\mathrm{ridge}}(t,x)$ and $U_{\mathrm{ridge}}(t,x)$ stay bounded as $t\uparrow T$;
  3. (3)
    the singularity is localized at the ridge apex $(x,\xi)=(0,\pm 1)$, in the sense that
    $V_{\mathrm{ridge}}(t,0)=\frac{6}{T-t}\to+\infty\qquad\text{as }t\uparrow T,$
    while
    $\frac{U_{\mathrm{ridge}}(t,x)}{U_{\mathrm{ridge}}(0,x)}=\frac{72}{2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}}$
    blows up like $(T-t)^{-2}$ at $x=0$.
In particular, the reduced ridge system (2.13) admits an explicit finite-time blow-up profile concentrated at the ridge apex.
Proof.
Substitute $(a,b)=(a_{\mathrm{ridge}}(x),b_{\mathrm{ridge}}(x))$ into (2.17). Since
$a_{\mathrm{ridge}}(0)=A,\qquad b_{\mathrm{ridge}}(0)=0,$
at $x=0$ we obtain
$V_{\mathrm{ridge}}(t,0)=-\frac{6\cdot 2A(tA-6)}{2(tA-6)^{2}}=\frac{6A}{6-tA}=\frac{6}{T-t},$
which blows up at $t=T=6/A$. Likewise,
$\frac{U_{\mathrm{ridge}}(t,x)}{U_{\mathrm{ridge}}(0,x)}=\frac{72}{2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}},$
and at $x=0$ this equals
$\frac{36}{(tA-6)^{2}}=\frac{T^{2}}{(T-t)^{2}},$
so the normalized $U$-profile also blows up at time $T$.
Now fix $x\neq 0$. Then $a_{\mathrm{ridge}}(x)<A$, hence
$Ta_{\mathrm{ridge}}(x)-6=\frac{6a_{\mathrm{ridge}}(x)}{A}-6<0.$
Therefore the denominator
$2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}$
stays strictly positive up to $t=T$, and both $V_{\mathrm{ridge}}(t,x)$ and $U_{\mathrm{ridge}}(t,x)$ remain bounded there. This proves the claimed localization of the singularity. ∎
The preceding theorem gives an explicit blowing-up ridge core. The full background problem is to extend that ridge core to a classical wedge solution of (2.9). The next proposition isolates the exact compatibility requirement behind this extension.
Theorem 2.7 (Compatibility condition for preserving ridge flatness).
Let $(U,V,G,P)$ be a classical solution of (2.9) on $0,t_{*}]\times\Omega$,
$\Omega:=\{(x,\xi):x>0,\ |\xi|\leq 1\},$
and fix a ridge $\xi_{0}\in\{\pm 1\}$. Assume that along $\xi=\xi_{0}$ one has
$G=V,\qquad U_{\xi}=V_{\xi}=G_{\xi}=P_{\xi}=0.$ (2.20)
Then the first normal derivative of $U$ evolves by
$\lambda\,\partial_{t}(U_{\xi})(t,x,\xi_{0})=-\xi_{0}V(t,x,\xi_{0})\Bigl(xU_{x}(t,x,\xi_{0})+2U_{\xi\xi}(t,x,\xi_{0})\Bigr).$ (2.21)
Consequently, exact first-order ridge flatness is not automatically propagated by (2.9). In particular, for the seed family
$\left\{\begin{aligned} &r=x^{2}+A_{1}(\xi^{2}-1)^{2m},\\ &a(x,\xi)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b(x,\xi)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0,\end{aligned}\right$. (2.22)
with $m\geq 2$, one has $U_{\xi\xi}(0,x,\xi_{0})=0$; hence (2.21) reduces initially to
$\lambda\,\partial_{t}(U_{\xi})(0,x,\xi_{0})=-\xi_{0}V(0,x,\xi_{0})\,xU_{x}(0,x,\xi_{0}),$
which is generically nonzero for $x\neq 0$.
Proof.
Differentiate the $U$-equation in (2.9) with respect to $\xi$. Writing
$A(t,x,\xi):=\frac{\xi^{2}G-V}{1+\xi^{2}},\qquad B(t,x,\xi):=(G+V)\xi,$
the $U$-equation becomes
$\lambda U_{t}-\tfrac{1}{2}VU=-A\,xU_{x}-B\,U_{\xi}.$
Differentiating gives
$\lambda(U_{\xi})_{t}-\tfrac{1}{2}V_{\xi}U-\tfrac{1}{2}VU_{\xi}=-A_{\xi}xU_{x}-AxU_{x\xi}-B_{\xi}U_{\xi}-BU_{\xi\xi}.$
Now evaluate at $\xi=\xi_{0}$ under (2.20). Since $\xi_{0}^{2}=1$, $G=V$, and $G_{\xi}=V_{\xi}=0$ there, we have
$A(t,x,\xi_{0})=0,\qquad A_{\xi}(t,x,\xi_{0})=\xi_{0}V(t,x,\xi_{0}),\qquad B(t,x,\xi_{0})=2\xi_{0}V(t,x,\xi_{0}).$
Using also $U_{\xi}=V_{\xi}=0$ on the ridge, we obtain exactly (2.21). Finally, for the seed family (2.22) with $m\geq 2$, the dependence on $\xi$ enters through $(\xi^{2}-1)^{2m}$, so the initial second $\xi$-derivative vanishes at $\xi=\pm 1$. The last assertion follows. ∎
Remark 2.8 (What revision (c) actually proves).
Revision (c) establishes an explicit blowing-up ridge background and pinpoints the compatibility condition needed to extend it to a full wedge background. The ridge core itself is completely explicit by Theorem 2.6. What remains open is the full-wedge extension problem: one must construct a classical solution of (2.9) on $0,T)\times\Omega$ whose ridge trace agrees with that explicit core and whose normal derivatives satisfy the compatibility relation (2.21) (and the analogous relations for $V,G,P$).
In particular, the previous heuristic claim that a sufficiently large power $m$ in (2.22) should by itself preserve exact ridge flatness is too optimistic. Large $m$ still improves angular flatness of the initial seed and is useful for weighted estimates, but Proposition 2.7 shows that exact flatness is not an invariant of the full background flow without additional structure. This is now the main analytic issue on the background side of the manuscript.
3. Perturbation PDEs
We derive the remainder system around the explicit background. The full solutions are written as the sum of the background fields and the remainders:
[…]
After substituting  (3.1) into  (2.6)(1,2,3,4), we separate the full system into the background equations for $(V,U,G,P)$ and the exact remainder equations for $(v,u,g,p)$. In the present revision all pure-background terms are kept in the background equations, so the remainder system contains only linear couplings to the background and genuinely nonlinear remainder–remainder interactions:
$\left\{\begin{aligned} \lambda u_{t}&=\tfrac{1}{2}(uV+Uv)-(g+v)\xi U_{\xi}-(G+V)\xi u_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xu_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xU_{x}+N_{1}\\[5.69054pt] \lambda v_{t}&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\qquad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{1}{\xi}p_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x}+N_{2}+\tfrac{1}{x}p_{x}\\[5.69054pt] \lambda g_{t}&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\quad\quad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{\mu}{\xi}p_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}+N_{3}+\tfrac{\mu}{x}p_{x}\\[5.69054pt] 0&=\xi xg_{x}+xv_{x}-({\xi^{2}+1})(g-v+\xi g_{\xi}+\xi v_{\xi})\end{aligned}\right$. (3.2)
Where the nonlinear perturbation terms are given by
$\left\{\begin{aligned} N_{1}:&=\tfrac{1}{2}uv\qquad\quad\,\,\,-(\xi^{2}g-v)xu_{x}-(g+v)\xi u_{\xi},\\[5.69054pt] N_{2}:&=v^{2}-u^{2}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xv_{x}-(g+v)\xi v_{\xi},\\[5.69054pt] N_{3}:&=-g^{2}\quad\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xg_{x}-(g+v)\xi g_{\xi}.\end{aligned}\right$. (3.3)
The divergence-free condition for the perturbation $(v,g)$ (3.2)(4) can be solved with the stream function $\psi(t,x,\xi)$ defined below:
[…]
3.1. Getting rid of $(p_{\xi},p_{x})$
Our next step is to get rid of $(p_{\xi},p_{x})$. Define $\omega$ as
[…]
So we obtain
[…]
Now we rewrite  (3.2)2,3 as:
[…]
where $(A,B)$ are defined as
$\left\{\begin{aligned} A:&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x},\\[5.69054pt] B:&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}.\end{aligned}\right$. (3.8)
The solutions for $(p_{x},p_{\xi})$ then become
[…]
Using $p_{x\xi}=p_{\xi x}$ to get rid of $p$, and using $g_{xt}$ of  (3.6) to simplify the result, we finally obtain:
$\left\{\begin{aligned} \tfrac{3}{2}\omega_{t}&=L_{2}+M_{2}\\[5.69054pt] L_{2}&=-\left(\tfrac{5}{3}xA_{x}+xB_{x}\right)-\left(1+\xi^{2}\right)\Bigl(\tfrac{5}{3\xi}A_{\xi}-\xi B_{\xi}\Bigr)\\[5.69054pt] M_{2}&=-\left(\tfrac{5}{3}xN_{2x}+xN_{3x}\right)-\left(1+\xi^{2}\right)\Bigl(\tfrac{5}{3\xi}N_{2\xi}-\xi N_{3\xi}\Bigr).\end{aligned}\right$. (3.10)
Substituting $(A,B)$ of  (3.8) into  (3.10)(2) and using $(g_{x},g_{xx},g_{x\xi},G_{x},G_{xx},G_{x\xi})$ of  (3.6) to simplify the results, and using (3.4) express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, we obtain
[…]
Substituting $N_{2}$, $N_{3}$ of  (3.3) into  (3.10)(3) and $(g_{x},g_{xx},g_{x\xi},G_{x},G_{xx},G_{x\xi})$ of  (3.6) to simplify the results, and using (3.4) express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, we obtain
$\left\{\begin{aligned} M_{2}&=\tfrac{10}{3\xi}\left(\left(\xi^{2}+1\right)u_{\xi}+\xi xu_{x}\right)u\\ &+\tfrac{\xi^{2}-1}{\xi^{2}+1}\left(x\psi_{x}+2\psi\right)\omega\\ &-\xi\left(x\psi_{x}+2\psi\right)\omega_{\xi}\\ &+\big(\xi x\psi_{\xi}-\tfrac{\xi^{2}-1}{\xi^{2}+1}x\psi\bigr)\omega_{x}.\end{aligned}\right$. (3.12)
Using (3.4) to express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, the equation for $u_{t}$ in  (3.2)(1) can also be converted into desired form:
$\tfrac{3}{2}u_{t}$ $=L_{1}+M_{1},$ (3.13)
$\left\{\begin{aligned} L_{1}&=\tfrac{1}{2}V\,\,u-(G+V)\xi u_{\xi}+\tfrac{1}{\xi^{2}+1}\left(V-\xi^{2}G\right)xu_{x}\\ &+\tfrac{1}{2}\left(U-4\xi U_{\xi}-\tfrac{2}{\xi^{2}+1}\left(\xi^{2}-1\right)xU_{x}\right)\psi\\ &+\tfrac{1}{2}\left(U+2xU_{x}\right)\xi\psi_{\xi}+\tfrac{1}{2}\left(\tfrac{\xi^{2}}{\xi^{2}+1}U-2\xi U_{\xi}\right)x\psi_{x}\end{aligned}\right$. (3.14)
$\left\{\begin{aligned} M_{1}&=\tfrac{1}{2}u\psi+\tfrac{1}{2}u\left(\xi\psi_{\xi}+\tfrac{\xi^{2}}{\xi^{2}+1}x\psi_{x}\right)\\ &-\tfrac{1}{\xi^{2}+1}\psi\left(2\left(\xi^{2}+1\right)\xi u_{\xi}+\left(\xi^{2}-1\right)xu_{x}\right)\\[5.69054pt] &+\left(xu_{x}\xi\psi_{\xi}-\xi u_{\xi}x\psi_{x}\right).\end{aligned}\right$. (3.15)
Substitution of  (3.4) into  (3.5) leads to
$\left\{\begin{aligned} \omega:&=\Delta\psi,\\ \Delta:&=-\tfrac{\left(5\xi^{2}+3\right)}{3\left(\xi^{2}+1\right)}x^{2}\psi_{xx}-\tfrac{\left(19\xi^{2}+21\right)}{3\left(\xi^{2}+1\right)}x\psi_{x}-\tfrac{4}{3}x\xi\psi_{x\xi}\\ &-\tfrac{10\left(\xi^{2}+1\right)}{3\xi}\psi_{\xi}-\tfrac{1}{3}\left(\xi^{2}+1\right)\left(3\xi^{2}+5\right)\psi_{\xi\xi}\end{aligned}\right$. (3.16)
4. Explicit background, perturbation PDEs, and energy bounds
4.1. Background Solutions
The background solutions $V(t,x,\theta),U(t,x,\theta)$ are updated with the new seed profiles from (4.2):
[…]
Initial conditions
$r:=x^{2}\bigl(1+\cos^{2}(2\theta)\bigr)\geq 0,\qquad V_{0}(x,\theta)=\frac{A}{1+r^{3}},\qquad U_{0}(x,\theta)=\frac{B\,r}{\bigl(1+r^{3}\bigr)^{2}}.$ (4.2)
4.2. Initial energy and finiteness
We record the “initial energy” (at $t=0$) associated with the background profile:
$E(0):=\int_{-\pi/4}^{\pi/4}\int_{-\infty}^{\infty}x^{2}\Big[(V_{0}(x,\theta))^{2}+U_{0}^{2}(x,\theta)\Big]\ |x|\,dx\,d\theta.$ (4.3)
Remark 4.1 (Behavior of the weighted energy for the new seed choice).
In the modified background model (with $U^{2}$ in the VV–equation) we take the energy density $V^{2}+U^{2}$. For the new seeds (4.2),
$r:=x^{2}\bigl(1+\cos^{2}(2\theta)\bigr),\qquad V_{0}(x,\theta)=\frac{A}{1+r^{3}},\qquad U_{0}(x,\theta)=\frac{B\,r}{(1+r^{3})^{2}},$
one still has $V_{0}(0,\theta)=A$ and $U_{0}(0,\theta)=0$ for every $\theta\in[-\pi/4,\pi/4]$, so the blow-up time remains $T=6/A$ and the singularity remains localized at the origin $x=0$. A Taylor expansion at $r=0$ gives
$V_{0}(x,\theta)=A\bigl(1-r^{3}+O(r^{6})\bigr),\qquad U_{0}(x,\theta)=B\,r+O(r^{4}),$
while at spatial infinity one has $r\sim x^{2}$, hence
$V_{0}(x,\theta)=O(x^{-6}),\qquad U_{0}(x,\theta)=O(x^{-10}).$
Therefore the initial weighted energy $E(0)$ is finite. The endpoint behavior at $t=T$ is different, however: the new seeds keep the background explicit and smooth for every fixed $t<T$, but they no longer produce a uniformly bounded endpoint energy at $t=T$.
4.3. Updated Linear PDEs and Nonlinear Terms
The final remainder system for $(u,\omega,\psi)$ is:
[…]
4.4. Coefficient Functions
The effective elliptic operator $\Delta$ in (4.4)(3) is defined by:
$\left\{\begin{aligned} \Delta&=c_{1}(\xi)x^{2}\psi_{xx}+c_{2}(\xi)x\psi_{x}+c_{4}(\xi)x\psi_{x\xi}+c_{3}(\xi)\psi_{\xi}+c_{5}(\xi)\psi_{\xi\xi}\\ c_{1}&=-\tfrac{\left(5\xi^{2}+3\right)}{3\left(\xi^{2}+1\right)},\quad c_{2}=-\tfrac{\left(19\xi^{2}+21\right)}{3\left(\xi^{2}+1\right)}\quad c_{4}=-\tfrac{4}{3}\xi\\ c_{3}&=-\tfrac{10}{3\xi}\left(\xi^{2}+1\right)\quad c_{5}=-\tfrac{1}{3}\left(\xi^{2}+1\right)\left(3\xi^{2}+5\right).\end{aligned}\right$. (4.5)
Remark 4.2.
The apparent singular factor $\tfrac{1}{\xi}$ in $c_{3}(\xi)$ is absorbed by the weighted derivative $D_{\xi}=\tfrac{1}{\xi}\partial_{\xi}$: writing
$c_{3}(\xi)=\widetilde{c}_{3}(\xi)\,\tfrac{1}{\xi},\qquad\widetilde{c}_{3}(\theta):=-\tfrac{10}{3}\left(\xi^{2}+1\right),$
we have $c_{3}(\xi)\,\partial_{\xi}=\widetilde{c}_{3}(\xi)\,D_{\xi}$ with bounded smooth $\widetilde{c}_{3}$ on [1,1][-1,1]. Likewise, terms involving $\tfrac{1}{\xi}$ are treated as bounded multipliers in the weighted energy once expressed in terms of $D_{\xi}$ (or placed into divergence form in $\xi$).
The coefficients $c_{1}(\xi),c_{2}(\xi),c_{4}(\xi),c_{5}(\xi)$ are bounded on [1,1][-1,1]. The coefficient $c_{3}(\xi)$ contains the factor $\tfrac{1}{\xi}$ and is therefore singular at $\xi=0$ in its raw form; however, as explained above, the combination $c_{3}(\xi)\partial_{\xi}$ is naturally rewritten as a bounded multiplier times the adapted derivative $D_{\xi}=\tfrac{1}{\xi}\partial_{\xi}$. This is the form used throughout the energy estimates.
4.5. Initial conditions and boundary conditions
Boundary conditions (for the $\xi$-edges and for $x\to\pm\infty$) and initial conditions are:
[…]
The phrase “sufficiently fast decay” as $\xi\to\pm 1$ means enough vanishing and regularity near the edge so that the weighted Sobolev norms used below are finite and the boundary terms produced by integration by parts vanish at $\xi=\pm 1$.
Remark 4.3.
If we define
$\left\{\begin{aligned} &y=\log x,\quad\Delta_{1}=\Delta\\ &\omega_{1}(t,y,\xi)=\omega(t,x,\xi),\\ &\psi_{1}(t,y,\xi)=\psi(t,x,\xi),\end{aligned}\right$. (4.7)
Then $\omega=\Delta\psi$ in (4.4) becomes
[…]
Thus (4.8) becomes a well-defined elliptic problem in the strip $\Omega=\{(y,\xi):y\in\mathbb{R},\ \xi\in[-1,1]\}$.
5. Linear estimates and conditional nonlinear control up to blow-up time
Updated perturbation system. Throughout this section we work with the final perturbation equations derived in Section 2. We study the perturbation system (3.13) and (3.10), together with the coefficient collections (3.14),(3.15),(3.11),(3.12) and the elliptic operator (4.5), on the time interval $[0,T)$ up to the background blow-up time, around the explicit background (4.1). Throughout, all Lebesgue and Sobolev norms are taken with respect to the weighted measure $d\mu_{w}=w(\xi)|x|\,dx\,d\xi$, and we use the desingularized angular derivative
$D_{\xi}:=\tfrac{1}{\xi}\partial_{\xi}.$
This is the natural derivative after the change of variables $\xi=\tan\theta$, which turns the trigonometric coefficients into rational functions of $\xi$.
5.1. Bootstrap framework and updated background coefficient bounds
Fix an integer $k\geq 6$. Define the perturbation energy
$\mathcal{E}_{k}(t):=\sum_{\begin{subarray}{c}j+\ell\leq k\end{subarray}}\Big(\|\partial_{x}^{j}D_{\xi}^{\ell}u(t)\|_{L^{2}_{\mu_{w}}}^{2}+\|\partial_{x}^{j}D_{\xi}^{\ell}\omega(t)\|_{L^{2}_{\mu_{w}}}^{2}\Big)+\sum_{\begin{subarray}{c}j+\ell\leq k+1\end{subarray}}\|\partial_{x}^{j}D_{\xi}^{\ell}\psi(t)\|_{L^{2}_{\mu_{w}}}^{2}.$ (5.1)
Background coefficient bounds actually needed in the energy method. The explicit background (4.1) provides $(U,V)$ in closed form. What the stability estimates require is not a uniform bound on the raw derivatives $\partial_{x}^{m}\partial_{\theta}^{\ell}V$ (which can grow faster than $(T-t)^{-1}$ near the intermediate scale $r^{2}\sim T-t$), but rather uniform control of the degenerate combinations that appear in (3.14),(3.11) and in the weighted Sobolev norms.
Define the adapted derivatives
$Z_{x}:=x\partial_{x},\qquad D_{\xi}:=\tfrac{1}{\xi}\partial_{\xi},$
Then for each integer $k\geq 0$ there exists $C_{*}=C_{*}(A,B,\text{seeds},k)$ such that for all $t\in[0,T)$ the following estimate holds.
Lemma 5.1 (Adapted background coefficient bounds).
$\sum_{j+\ell\leq k}\Big(\|Z_{x}^{j}D_{\xi}^{\ell}V(t)\|_{L^{\infty}}+\|Z_{x}^{j}D_{\xi}^{\ell}U(t)\|_{L^{\infty}}\Big)$ (5.2)
$+\sum_{j+\ell\leq k}\Big(\Big\|Z_{x}^{j}D_{\xi}^{\ell}\bigl(x\,V_{x}(t)\bigr)\Big\|_{L^{\infty}}+\Big\|Z_{x}^{j}D_{\xi}^{\ell}\bigl(x\,U_{x}(t)\bigr)\Big\|_{L^{\infty}}\Big)$
$\;\leq\;\frac{C_{*}}{T-t}.$
Proof sketch.
The estimate is a direct consequence of the closed-form background formula together with the seed expansions near the ridge apex and the fact that away from the apex the denominator in (4.1) is bounded below uniformly. Near the apex one uses the denominator scale $D(t,x,\theta)\gtrsim(T-t)^{2}+r^{2}$, together with the observation that $Z_{x}=x\partial_{x}$ acts like $r\partial_{r}$ and $D_{\xi}$ neutralizes the harmless trigonometric singular factors in angular derivatives. ∎
In particular,
$\|V(t)\|_{L^{\infty}}=\frac{6}{T-t}\quad\text{(attained at }(x,\xi)=(0,\pm 1)\text{)},\qquad\|U(t)\|_{L^{\infty}}\lesssim\frac{1}{T-t}.$
5.2. Elliptic control of $\psi$ from $\omega=\Delta\psi$
The elliptic relation in (4.4) reads $\omega=\Delta\psi$, where $\Delta$ is given by (4.5). After rewriting the angular part in terms of the adapted derivative $D_{\xi}$ (as already indicated in the weighted-norms subsection), the operator $\Delta$ has the same principal structure as $\Delta_{1}=\partial_{\log x}^{2}+D_{\xi}^{2}$ with lower-order $\xi$-dependent coefficients controlled on the wedge. Accordingly, we record the following weighted elliptic estimate as the analytic input needed for the perturbation argument: for all integers $m\geq 0$,
$\|\psi(t)\|_{H^{m+2}_{\mu_{w}}}\leq C_{\Delta,m}\,\|\omega(t)\|_{H^{m}_{\mu_{w}}},$ (5.3)
where the constant depends only on the wedge geometry, the boundary conditions, and the bounded coefficient functions appearing in (4.5). This estimate is natural from the $y=\log x$ reformulation discussed in Remark 4.3; in the present manuscript we use it as a working elliptic input for the $\psi$-estimate rather than as a separately proved theorem.
In particular, since $k\geq 6$, Sobolev embedding in the $(x,\xi)$ variables (with $D_{\xi}$ counted as one derivative) gives
$\|u(t)\|_{L^{\infty}}+\|\omega(t)\|_{L^{\infty}}+\|\psi(t)\|_{W^{1,\infty}}\leq C\,\mathcal{E}_{k}(t)^{1/2}.$ (5.4)
5.3. Energy inequality for $(u,\omega)$
Differentiate the $u$-equation and the $\omega$-equation in (4.4) by $\partial_{x}^{j}D_{\xi}^{\ell}$ for $j+\ell\leq k$, take the $L^{2}_{\mu_{w}}$ inner product with $\partial_{x}^{j}D_{\xi}^{\ell}u$ and $\partial_{x}^{j}D_{\xi}^{\ell}\omega$, and sum over $j+\ell\leq k$. The transport terms are now written directly in the $(x,\xi)$ variables, so the integration-by-parts step is carried out in $x$ and $\xi$. The boundary contributions vanish because of the remainder boundary conditions at $\xi=\pm 1$, the decay as $|x|\to\infty$, and the weighted formulation using $D_{\xi}=\xi^{-1}\partial_{\xi}$.
Using the commutator estimates and (5.4), one obtains an inequality of the form
$\frac{d}{dt}\mathcal{E}_{k}(t)\leq\frac{C_{\rm lin}}{T-t}\,\mathcal{E}_{k}(t)+C_{\rm nl}\,\Big(\|M_{1}(t)\|_{H^{k}_{\mu_{w}}}+\|M_{2}(t)\|_{H^{k}_{\mu_{w}}}\Big)\,\mathcal{E}_{k}(t)^{1/2}.$ (5.5)
Quadratic remainder terms. From the explicit forms of $M_{1},M_{2}$ in (3.15) and (3.12), together with Moser and Sobolev product estimates in the $(x,\xi)$ variables, one obtains
$\|M_{1}(t)\|_{H^{k}_{\mu_{w}}}+\|M_{2}(t)\|_{H^{k}_{\mu_{w}}}\leq C\,\mathcal{E}_{k}(t).$
Because all pure-background terms have been kept in the background system, there is no additive forcing term in the remainder energy inequality. Thus it is natural to rewrite (5.5) in terms of
$Y(t):=\mathcal{E}_{k}(t)^{1/2}.$
Then
$Y^{\prime}(t)\leq\frac{C_{\rm lin}}{T-t}Y(t)+C_{\rm nl}Y(t)^{2}$ (5.6)
whenever $Y(t)>0$.
The important point is that (5.6) by itself does not yet imply a closed bootstrap with a remainder strictly smaller than the background singularity. What it does give is an Elgindi-type conditional transfer principle: if the remainder stays in a class whose growth is weaker than the background blow-up rate, then the quadratic term is perturbative and the background singularity transfers to the full solution.
To make this precise, fix an exponent $\sigma>0$ and define the renormalized energy envelope
$X_{\sigma}(t):=(T-t)^{\sigma}Y(t).$
Differentiating and using (5.6) gives
$X_{\sigma}^{\prime}(t)\leq\frac{C_{\rm lin}-\sigma}{T-t}X_{\sigma}(t)+C_{\rm nl}(T-t)^{-\sigma}X_{\sigma}(t)^{2}.$ (5.7)
Hence, whenever
$\sigma>C_{\rm lin},$ (5.8)
and whenever a bootstrap bound of the form
$X_{\sigma}(t)\leq M\varepsilon\qquad\text{for }0\leq t\leq t_{*}$ (5.9)
holds with $\varepsilon>0$ sufficiently small, the right-hand side of (5.7) is integrable and the quadratic term can be absorbed. Standard continuity then yields
$X_{\sigma}(t)\leq 2X_{\sigma}(0)\qquad\text{for }0\leq t\leq t_{*}.$ (5.10)
Equivalently,
$Y(t)\leq 2Y(0)\Bigl(\frac{T}{T-t}\Bigr)^{\sigma},\qquad 0\leq t\leq t_{*}.$ (5.11)
In particular, if one can choose $\sigma<1$ while still having (5.8), then the remainder stays strictly below the background blow-up scale $(T-t)^{-1}$ in the detecting norm.
This discussion is summarized in the following conditional theorem.
Theorem 5.2 (Conditional nonlinear control up to the background blow-up time).
Assume that the explicit background solution exists on $[0,T)$, blows up at $(x,\xi)=(0,\pm 1)$ at time $T=6/A$, and satisfies the adapted coefficient bounds of Lemma 5.1. Assume also that the weighted elliptic estimate (5.3) holds. Let $k\geq 6$, and let $(u,\omega,\psi)$ solve the exact remainder system on $0,t_{*}]\subset[0,T)$.
Then there exist constants $C_{\rm lin},C_{\rm nl}>0$, depending only on $k$ and the background coefficient bounds, such that (5.6) holds. Consequently, for every exponent $\sigma$ satisfying (5.8), there exists $\varepsilon_{0}=\varepsilon_{0}(\sigma,k)>0$ with the following property: if
$X_{\sigma}(0)=T^{\sigma}\mathcal{E}_{k}(0)^{1/2}\leq\varepsilon_{0},$
and if the bootstrap assumption (5.9) holds on $0,t_{*}]$, then in fact (5.10) holds on $0,t_{*}]$.
Remark 5.3 (What this proves now, and what still has to be improved).
Theorem 5.2 is already strong enough to put the remainder analysis into the same logical class as the Elgindi–Jeong mechanism: the singular core is the explicit background, and the nonlinear argument reduces to showing that the remainder remains in a better class. However, the theorem is still conditional. To turn it into a full stability statement one still needs an independent argument guaranteeing a gap (5.8) with some exponent $\sigma<1$ in the norm that detects the background blow-up. This may come from sharper coercivity, additional vanishing of the remainders at the ridge, or a more scale-adapted energy functional.
Under this conditional control, one obtains a blow-up transfer statement for the full solution.
Theorem 5.4 (Conditional transfer of background blow-up to the full solution).
Assume the hypotheses of Theorem 5.2. In addition, suppose that the chosen detecting norm $\mathcal{N}_{\rm det}(t)$ for the full solution satisfies
$\mathcal{N}_{\rm det}^{\rm bg}(t)\sim c_{0}(T-t)^{-1}\qquad\text{for some }c_{0}>0,$
when evaluated on the background, and that the remainder contribution is estimated by
$\mathcal{N}_{\rm det}^{\rm rem}(t)\leq C_{\rm det}Y(t)$
for $0\leq t<T$. If there exists $\sigma\in(C_{\rm lin},1)$ such that (5.10) holds on $[0,T)$, then
$\mathcal{N}_{\rm det}(t)=\mathcal{N}_{\rm det}^{\rm bg}(t)+O\bigl((T-t)^{-\sigma}\bigr)\qquad\text{as }t\uparrow T,$
and hence the full solution blows up at time $T$ with the same leading-order singularity location and blow-up scale as the background.
Accordingly, the logical bottleneck of the manuscript is no longer a forcing obstruction in the remainder equations. The main unresolved issue is instead the rigorous construction of the background solution up to its blow-up time with the required ridge flatness and coefficient bounds, together with whatever refined estimate is needed to produce a genuine gap exponent $\sigma<1$ in the remainder norm. Once those two inputs are available, the present stability mechanism upgrades directly to a nonlinear remainder theorem in the spirit of Elgindi.
6. Conclusion
We derived a closed (1+2)(1+2)D subsystem (E1) from the inviscid 2D Boussinesq equations under a parity ansatz and organized its blow-up analysis around two components: a singular background core and an exact remainder system. In the updated formulation, the ridge rays carry a convection-free (1+1)(1+1)D CLM-type reaction dynamics with explicit finite-time blow-up, while the full wedge problem is rewritten in the $(x,\xi)$ variables so that the remainder equations are exact and all pure-background terms remain in the background system.
The main gain of this revision is conceptual and technical at the same time. On the remainder side, the earlier large forcing obstruction has disappeared. What remains is a singular linear problem with quadratic lower-order interactions, and this is precisely the setting in which an Elgindi-type transfer mechanism becomes available. The weighted energy method developed in Section 5 now shows that, if the background exists on $[0,T)$ with the coefficient bounds required there, and if the remainder stays subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
Accordingly, the main unresolved step is now sharply identified. It is no longer the derivation of the remainder system, nor the linear weighted estimate, but the full-wedge background construction up to the blow-up time together with the rigidity properties needed to propagate the ridge structure and close the nonlinear bootstrap without loss. Revision (c) makes this explicit: the ridge blow-up mechanism itself is available in closed form, but extending it to a full background with the necessary compatibility bounds remains the decisive open problem.
This sharpened formulation is useful even before the final nonlinear theorem is complete. It separates what is already rigorous from what is still conditional, and it places the manuscript much closer in spirit to the singular-core-plus-remainder logic used by Elgindi and collaborators. At the present stage, the paper provides an exact derivation from Boussinesq, an explicit ridge blow-up mechanism, an exact remainder system, singular weighted linear control, and a precise conditional blow-up transfer statement.
Natural next steps are therefore clear. The first is to prove the full background existence and blow-up theorem for the anisotropic seed family used here. The second is to sharpen the detecting norm so that the remainder remains strictly below the background blow-up rate, yielding a closed nonlinear bootstrap. After that, one can revisit modulation of geometric parameters and lower-regularity weighted theories.
7. Acknowledgements
ChatGPT is credited as a substantive contributor to drafting and technical editing; responsibility for correctness remains with the author. The author thanks Prof. Zixiang Zhou of the Department of Mathematics at Fudan University and Prof. Jie Qin of the Department of Mathematics at the University of California, Santa Cruz for their continuous support and encouragement over the years. The author also thanks colleagues and the broader PDE and fluid-dynamics community for stimulating discussions on inviscid Boussinesq and CLM-type models. (Computational assistance: ChatGPT, GPT–5.2 Thinking, OpenAI; sessions in March 2026.)