### 03 Apr 2026 v3

# A pressure–velocity apex blow-up mechanism and conditional perturbative stability framework for a (1+2)(1+2)D system (E1) derived from the 2D inviscid Boussinesq equations

Abstract.
In polar variables on a planar wedge, we study a closed (1+2)(1+2)D system (E1) derived from the two-dimensional inviscid Boussinesq equations under a parity ansatz. A central feature of the paper is the use of the velocity–pressure formulation: it keeps the divergence-free structure visible, reveals the flat ridge-ray geometry at least on short time intervals, and leads to an exact apex-dynamics reduction on the distinguished ridge rays. This reduction produces an explicit finite-time blow-up mechanism at the ridge apex $(x,\xi)=(0,\pm 1)$.
The paper has three main outputs. First, we identify the special ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting exact apex blow-up mechanism for $t\in[0,T]$. Second, we derive the exact remainder equations around a prescribed background and prove singular weighted linear estimates for the remainder system. Third, we formulate a conditional nonlinear control principle in the spirit of Elgindi–Jeong: if a compatible background exists on $[0,T)$ with the coefficient bounds required by the weighted energy method, and if the remainder remains subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
Our approach is complementary to the vorticity–stream and boundary-driven singularity frameworks in the recent literature. Here the analysis is carried out on the full reduced-plane geometry attached to the pressure–velocity reduction, with smooth functions and with symmetry replacing boundary/irregularity mechanisms in the handling of convection. What is unconditional in the present paper is the exact reduction from Boussinesq, the exact apex blow-up dynamics on the ridge, the apex flatness criterion at $x=0$, and the weighted remainder framework. What remains conditional is the construction of a full background away from the apex together with the rigidity properties needed to close the bootstrap without loss. In this sense, the manuscript isolates the background-extension problem as the main remaining step toward a complete nonlinear stability theorem for the blow-up scenario.
Key words and phrases:
2D inviscid Boussinesq, Constantin–Lax–Majda type model, finite-time blow-up, explicit solutions, conditional perturbative framework, ridge reduction, weighted Sobolev energy, divergence-form closure
2020 Mathematics Subject Classification:
35B44, 35B40, 35Q86, 76B03, 76D05
2020 Mathematics Subject Classification. Primary 35B44, 35Q86, 76B03; Secondary 35B35, 35Q35.
Keywords. 2D inviscid Boussinesq equations; Constantin–Lax–Majda type model; explicit blow-up solutions; ridge reduction; weighted energy; linear estimates; conditional nonlinear control; weighted Sobolev energies.
1. Introduction
The question of whether smooth solutions can develop a singularity in finite time is a central theme in nonlinear PDE and fluid mechanics. In this paper we study a $(1+2)$-dimensional closed subsystem (E1), rigorously derived from the inviscid 2D Boussinesq equations in velocity–pressure form under a parity ansatz. Our aim is twofold: first, to identify an explicit finite-time blow-up mechanism at the ridge apex; second, to formulate a perturbative stability theory around compatible backgrounds that is mathematically solid at the linear level and explicit about the remaining nonlinear obstruction.
A central advantage of the pressure–velocity formulation is that the divergence-free condition remains visible throughout the reduction, the flat ridge-ray structure is revealed directly at least for short times, and the apex dynamics on the distinguished rays can be isolated exactly on the full reduced-plane geometry. This point of view is complementary to formulations based on vorticity–stream variables. In the present paper, convection control is handled through the inherited symmetry of the reduced equations rather than through a physical boundary or a lower-regularity functional class. The resulting framework is therefore best described as a complementary pressure–velocity and smooth-function approach to finite-time blow-up for structures derived from the Boussinesq system.
We will call the Hou-Li type new variables $\{u,v,g\}$ of  (2.3) the building blocks of vorticity, because their units are equal to the unit of $\boldsymbol{\omega}=\nabla\times\boldsymbol{u}$. Also the quadratic vortex stretching terms are greatly simplified: $(uv,\,v^{2}-u^{2},\,-g^{2})$.
Related work and context.
As emphasized by Elgindi–Jeong [17] and Elgindi–Pasqualotto [19], singularity formation for the inviscid 2D Boussinesq and 3D Euler systems has a long history. We recall only a small selection of representative references here, emphasizing works most closely aligned with the blow-up mechanism and stability framework developed below. For the inviscid 2D Boussinesq system, classical local theory and conditional breakdown criteria go back to Cannon–DiBenedetto [12], Chae–Nam [3] (see also Chae–Kim–Nam [7]), and Taniuchi [29]; see also Wu’s lecture notes [30] and the Euler perspective of Beale–Kato–Majda [1] and Constantin [10, 11]. On the modeling side, explicit/didactic mechanisms include Chae–Constantin–Wu [2] and the CLM/De Gregorio lineage [9, 16], as well as 1D reductions for Boussinesq-type dynamics such as Choi–Kiselev–Yao [8]. For rigorous singularity constructions and perturbative stability scenarios in PDE settings, see Elgindi–Jeong [18, 17], Chen–Hou [4, 5], and the synthesis of Drivas–Elgindi [15]. In comparison with those works, the present paper adopts a different viewpoint: it starts from the pressure–velocity form, works with smooth functions on the full reduced-plane geometry associated with the symmetry reduction, and exploits symmetry rather than boundary effects or lower-regularity singular norms in the handling of convection. The benefit of this viewpoint is that the divergence-free structure and the ridge-ray geometry remain directly visible in the reduced equations, which in turn makes the exact apex-dynamics reduction transparent.
Main achievements.
  • (0)
    We derive the closed subsystem (E1) exactly from the inviscid 2D Boussinesq equations under a parity ansatz and identify the variables $\{u,v,g\}$ as convenient vorticity building blocks.
  • (1)
    We identify the ridge rays on which (E1) reduces to a convection-free (1+1)(1+1)D reaction system of Constantin–Lax–Majda type, and we record the resulting exact finite-time blow-up profile at the ridge apex $(x,\xi)=(0,\pm 1)$ for $t\in[0,T]$.
  • (2)
    We show that the pressure–velocity form reveals the divergence-free structure and the short-time flat ridge-ray geometry in a way that is compatible with the exact apex reduction.
  • (3)
    We derive the exact remainder equations around a prescribed background in the $(x,\xi)$ variables, with all pure-background contributions retained in the background system.
  • (4)
    We prove weighted singular linear estimates and formulate a conditional nonlinear remainder theorem of Elgindi type: once a compatible background is available with the required coefficient bounds, blow-up transfers from its apex dynamics to the full solution.
  • (5)
    We isolate the remaining open step in the program, namely the construction and control of a full background away from the apex together with the compatibility structure needed to close the nonlinear bootstrap.
Organization. Section 2 derives the closed subsystem (2.6) from the inviscid 2D Boussinesq equations and identifies the ridge rays on which (E1) reduces to a (1+1)(1+1)D CLM-type reaction system. Section 2.6 rewrites the background and remainder system in the $(x,\xi)$ variables. Section 4.4 records the exact remainder equations, and Section 4 collects the candidate background formulas, the explicit apex dynamics, and the corresponding weighted pointwise and energy bounds used in the conditional framework. Section 5 proves the singular weighted linear estimates, formulates the conditional nonlinear remainder mechanism, and explains how blow-up transfers from the background to the full solution once the remaining compatibility and background hypotheses are verified.
2. The derivation of system (E1) from 2D inviscid Boussinesq equations
2.1. The derivation of system (E1) from 2D inviscid Boussinesq equations
In this section we rewrite the velocity–pressure form of the 2D inviscid Boussinesq equations (see, for example, Wu [30], Elgindi–Jeong [18], and Kiselev–Pu–Yao [24]) in terms of new variables that we call vorticity building blocks. In these variables, the stretching and transport structure becomes more transparent.
In the velocity-pressure form, the 2D inviscid Boussinesq equations for velocity $\boldsymbol{u}=u_{2}\boldsymbol{e}_{2}+u_{3}\boldsymbol{e}_{3}$, pressure $P$ and density $\vartheta$ (assuming the gravity in $\boldsymbol{e}_{2}$ direction) in $\mathbb{R}^{2}$ are given by
[…]
$\partial_{2}^{2}P+\partial_{3}^{2}P=\partial_{2}\vartheta.$ (2.2)
Assume that $u_{2}$ is odd in $x_{2}$ and even in $x_{3}$, that $u_{3}$ is even in $x_{2}$ and odd in $x_{3}$, that $p$ is even in $(x_{2},x_{3})$, and that $\vartheta$ is odd in $x_{2}$ and even in $x_{3}$. Define the Hou–Li [23] type variables
$\{v,g,u^{2},p\}:=\left\{-\frac{u_{2}}{x_{2}},\frac{u_{3}}{x_{3}},\frac{\vartheta}{x_{2}},P\right\}$ (2.3)
Then  (2.1) can be converted into the following system (with $(\rho,z)=(x_{2},x_{3})$ for notational convenience):
Verification of the reduction. Under the stated parity assumptions, the whole plane system (2.1) can be extended to $\mathbb{R}^{2}$ by odd/even reflection. In particular, along the symmetry axes $x_{2}=0$ and $x_{3}=0$ the quotients
$v:=-\frac{u_{2}}{x_{2}},\qquad g:=\frac{u_{3}}{x_{3}},\qquad u^{2}:=\frac{\vartheta}{x_{2}}$ (2.4)
extend smoothly (the numerator has the matching odd symmetry), and we may write
$u_{2}=-\rho\,v,\qquad u_{3}=z\,g,\qquad\vartheta=\rho\,u^{2}.$
With this convention the material derivative becomes
$\tfrac{\mathrm{D}}{\mathrm{D}t}=\partial_{t}+u_{2}\partial_{\rho}+u_{3}\partial_{z}=\partial_{t}-\rho v\,\partial_{\rho}+zg\,\partial_{z},$ (2.5)
which matches the last line in (2.6). Now:
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}\vartheta=0$ and $\vartheta=\rho u^{2}$,
    $0=\tfrac{\mathrm{D}}{\mathrm{D}t}(\rho u^{2})=\Big(\tfrac{\mathrm{D}\rho}{\mathrm{D}t}\Big)u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=u_{2}u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=-\rho v\,u^{2}+\rho\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2}),$
    hence $\tfrac{\mathrm{D}}{\mathrm{D}t}(u^{2})=vu^{2}$, equivalently $\tfrac{\mathrm{D}}{\mathrm{D}t}u=\frac{1}{2}uv$.
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}u_{2}=-\partial_{\rho}P+\vartheta$ and $u_{2}=-\rho v$,
    $\tfrac{\mathrm{D}}{\mathrm{D}t}u_{2}=\tfrac{\mathrm{D}}{\mathrm{D}t}(-\rho v)=-(\tfrac{\mathrm{D}\rho}{\mathrm{D}t})v-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t}=-u_{2}v-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t}=\rho v^{2}-\rho\tfrac{\mathrm{D}v}{\mathrm{D}t},$
    so dividing by $\rho>0$ gives $\tfrac{\mathrm{D}v}{\mathrm{D}t}=v^{2}-u^{2}+\frac{1}{\rho}P_{\rho}$.
  • From $\tfrac{\tilde{\mathrm{D}}}{\mathrm{D}t}u_{3}=-\partial_{z}P$ and $u_{3}=zg$,
    $\tfrac{\mathrm{D}}{\mathrm{D}t}u_{3}=\tfrac{\mathrm{D}}{\mathrm{D}t}(zg)=(\tfrac{\mathrm{D}z}{\mathrm{D}t})g+z\tfrac{\mathrm{D}g}{\mathrm{D}t}=u_{3}g+z\tfrac{\mathrm{D}g}{\mathrm{D}t}=zg^{2}+z\tfrac{\mathrm{D}g}{\mathrm{D}t},$
    hence $\frac{\mathrm{D}g}{\mathrm{D}t}=-g^{2}-\frac{1}{z}P_{z}$.
  • Finally, incompressibility $\partial_{\rho}u_{2}+\partial_{z}u_{3}=0$ with $u_{2}=-\rho v$, $u_{3}=zg$ yields
    $0=\partial_{\rho}(-\rho v)+\partial_{z}(zg)=-(v+\rho v_{\rho})+(g+zg_{z}),$
    i.e. $zg_{z}-\rho v_{\rho}+g-v=0$.
This proves that the change of variables reduces (2.1) exactly to (2.6).
[…]
And the Poisson presure becomes
$P_{\rho\rho}+P_{zz}=u^{2}+\rho\,(u^{2})_{\rho}.$ (2.7)
Remark 2.1.
Inspection of  (2.6) shows that if the initial data for $\{u,v,g,p\}$ are symmetric in both $r$ and $z$, then the PDE system preserves these symmetries. In this sense,  (2.6) may be regarded as posed on $\mathbb{R}^{2}$.
Remark 2.2.
We will call $\{u,v,g\}$ the building blocks of vorticity, because their units are equal to the unit of $\boldsymbol{\omega}=\nabla\times\boldsymbol{u}$. Also the quadratic vortex stretching terms are greatly simplified: $(uv,\ v^{2}-u^{2},\ -g^{2})$.
2.2. System (E11)
We now introduce a stream function $\bar{\psi}$ and augment  (2.6) with two additional relations, obtaining system  (2.8). To reserve the symbols $(u,v,g,p)$ for later perturbation variables, we place bars on the background unknowns. Thus system (E11) in  (2.8) consists of five dependent variables $(\bar{u},\bar{v},\bar{g},\bar{p},\bar{\psi})$, viewed as even functions of $(\rho,z)$ on the meridian plane, together with six equations; the last equation defines $\tfrac{\mathrm{D}}{\mathrm{D}t}$.
$\text{(i) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{u}-\tfrac{1}{2}\,\bar{u}\,\bar{v},\quad(t,\rho,z)\in[0,T)\times\mathbb{R}^{2}$ (2.8a)
$\text{(ii) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{v}-\bar{v}^{2}+\bar{u}^{2}-\tfrac{1}{\rho}\bar{p}_{\rho},$ (2.8b)
$\text{(iii) }0$ $=\tfrac{\mathrm{D}}{\mathrm{D}t}\,\bar{g}+\bar{g}^{2}+\tfrac{{\mu^{2}}}{z}\,\,\bar{p}_{z},$ (2.8c)
$\text{(iv) }0$ $=z\partial_{z}\bar{g}-\rho\partial_{\rho}\bar{v}+\bar{g}-\bar{v},$ (2.8d)
$\text{(v) }0$ $=\bar{v}-\bar{\psi}-z\partial_{z}\bar{\psi},$ (2.8e)
$\text{(vi) }0$ $=\bar{g}-\bar{\psi}-\rho\partial_{\rho}\bar{\psi},$ (2.8f)
$\text{(vii) }\tfrac{\mathrm{D}}{\mathrm{D}t}:$ $={{\color[rgb]{0,0,1}\lambda}}\partial_{t}+\bar{g}\,z\partial_{z}-\bar{v}\,\rho\partial_{\rho},.$ (2.8g)
[…]
We will study the system in polar coordinates.
2.3. Polar coordinates $(x=r=\sqrt{\rho^{2}+z^{2}},\ \theta=\arctan(z/\rho))$
We use polar coordinates on the meridian plane:
$\rho=x\cos\left(\theta\right),\quad z=x\sin\left(\theta\right).$ (2.9)
Remark 2.4.
The polar coordinates $(x,\theta)$ on the meridian plane $(\rho,z)$ are also the spherical coordinates $(x,\theta,\phi)$ (with north pole at $\theta=\pi/2$) for 3D axisymmetric functions in $\mathbb{R}^{3}$.
2.4. Background
We write the background solutions as
$\bar{u}$ $=U(t,x,\xi),\quad\bar{v}=V(t,x,\xi),\quad\bar{g}=G(t,x,\xi),$ (2.10)
$\bar{p}$ $=P(t,x,\xi),\quad\xi=\tan(\theta).$
After substituting  (3.1) into  (2.8),we obtain four equations with the following structure:
$\left\{\begin{aligned} \lambda U_{t}-\tfrac{1}{2}V\,U&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xU_{x}-(G+V)\xi U_{\xi},\\[5.69054pt] \lambda V_{t}-V^{2}+U^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xV_{x}-(G+V)\xi V_{\xi}+\tfrac{1}{x}P_{x}-\tfrac{1+\xi^{2}}{x^{2}}\xi{P_{\xi}},\\[5.69054pt] \lambda G_{t}+G^{2}&=-\tfrac{1}{1+\xi^{2}}\left(\xi^{2}G-V\right)xG_{x}-\tfrac{1}{2}\xi(G^{2}+V^{2})_{\xi}-\tfrac{\mu}{x}P_{x}-\tfrac{\mu(1+\xi^{2})}{x^{2}\xi^{2}}\xi{P_{\xi}},\\[5.69054pt] G-V&=-\tfrac{1}{1+\xi^{2}}(\xi^{2}xG_{x}-xV_{x})-\xi(G_{\xi}+V_{\xi}).\end{aligned}\right$. (2.11)
The Poisson pressure equation (2.7) becomes
$\left\{\begin{aligned} \Delta P:&=P_{xx}+\tfrac{1}{x}P_{x}+\tfrac{2}{x^{2}}\left(\xi^{2}+1\right)\xi P_{\xi}+\tfrac{1}{x^{2}}\left(\xi^{2}+1\right)^{2}P_{\xi\xi}\\ &=U^{2}-\tfrac{2}{x}(\xi^{2}+1)^{1/2}\xi(U^{2})_{\xi}+(\xi^{2}+1)^{-1/2}(U^{2})_{x}.\end{aligned}\right$. (2.12)
We now examine how the equations simplify under the following Ansatz (I):
$\xi=\xi_{0}=\pm 1,\quad G(t,x,\xi_{0})=V(t,x,\xi_{0}).$ (2.13)
and ridge-flat Ansatz (II) (ridge-flatness condition in the directions normal to the ridge):
$(V_{\xi},U_{\xi},G_{\xi},P_{\xi})|_{\xi_{0}}=0,$ (2.14)
2.5. Ridge ray and Ridge functions
Theorem 2.5.
System  (2.8)(including the divergence-free condition  (2.8)(4)) restricted to the rays determined by $\xi^{2}=\xi_{0}^{2}=1$
Assuming Ansatz (I) and Ansatz (II), then we have:
(A): Divergence constraint and ridge flatness fixes the ridge rays. Under (2.13) and (2.14), the divergence identity (2.8d) implies
$\xi_{0}=\tan(\theta_{0})=\pm 1.$
Equivalently,
$\theta_{0}\in\Bigl\{\pm\tfrac{\pi}{4},\ \pi\pm\tfrac{\pi}{4}\Bigr\}.$
In the (r,z)(r,z)–plane this corresponds to the two straight lines through the origin
$z=\pm\,r,$
i.e. four rays (two rays on each diagonal line). When working in the first quadrant $r\geq 0$, $z\geq 0$, we take the principal choice
$\theta_{0}=\tfrac{\pi}{4}.$
(B). Convection-free on the principal ridge ray. On $\xi=\xi_{0}$,
$\left.\tfrac{\mathrm{D}}{\mathrm{D}t}\right|_{\xi=\xi_{0}}=\partial_{t}.$
(C) With the specific choice of scaling parameters $(\lambda,\mu)$:
$\lambda=\tfrac{3}{2},\quad\mu=\tfrac{5}{3},$ (2.15)
the dynamics of all ridge functions $\{U,V,G,P\}(t,x,\theta_{0})$ are completely determined by the following $1+1$-dimensional convection-free reaction system for $U(t,x,\xi_{0})$ and $V(t,x,\xi_{0})$.
$\left\{\begin{aligned} &U_{t}=\tfrac{1}{3}\,V\,U,\\[5.69054pt] &V_{t}=\tfrac{1}{6}V^{2}-\tfrac{5}{12}U^{2}.\end{aligned}\right$. (2.16)
and a simplified Poisson equation for $P(t,x,\xi_{0})$.
$P_{x}$ $=\tfrac{3x}{8}\left(U^{2}-2V^{2}\right),$ (2.17)
Notice that (2.16) is point-wise ODE system. It holds for every point $x\in\mathbb{R}$.
After setting $\tau=6t$, $\left(\tfrac{2}{5}\right)^{1/2}U(t,x)=\tfrac{1}{2}\omega(\tau,x)$, and $V(t,x)=\tfrac{1}{2}H\left(\omega(\tau,x)\right)$, we see that this system becomes algebraically equivalent, at each fixed $x$, to the Constantin–Lax–Majda reaction ODE system proposed as a model for vorticity stretching.
$\left\{\begin{aligned} &\partial_{\tau}\omega=\omega H(\omega)\\[5.69054pt] &\partial_{\tau}H(\omega)=\tfrac{1}{2}\left(H(\omega)^{2}-\omega^{2}\right).\end{aligned}\right$. (2.18)
Proof of theorem 2.5.
Claim (A) follows directly from  (3.2).
Because $\xi_{0}^{2}=1$ and $G(t,x,\xi_{0})=V(t,x,\xi_{0})$ on the ridge rays, the convection terms $(\xi^{2}G-V)\partial_{x}$ vanish in  (3.2). This proves Claim (B).
Thus our ridge ray system is composed of 4 PDEs, as shown below
$\left\{\begin{aligned} \tfrac{3}{2}U_{t}&=\tfrac{1}{2}V\,U,\\[5.69054pt] \tfrac{3}{2}V_{t}&=V^{2}-U^{2}+\tfrac{1}{x}P_{x},\\[5.69054pt] \tfrac{3}{2}V_{t}&=-V^{2}-\tfrac{5}{3x}P_{x}.\end{aligned}\right$. (2.19)
Separating $P_{x}$ from $V_{t}$ yields  (2.16) and eq:ridge-dynamics-2. This proves Claim (C) and completes the Proof of theorem 2.5. ∎
It is well known that system  (2.16), like the Constantin–Lax–Majda system, admits the following closed-form solutions.
[…]
Denote $\phi(\xi,\xi_{0})$ the famous end-vanishing smooth function on the inverval (refs needed here)
$\phi(\xi)=\exp\bigl(-(\xi^{2}-1)^{-2}\bigr),\quad\xi\in[-1,1].$ (2.21)
The function $\phi(\xi)$ has following nice properties:
[…]
Then we have the following:
Lemma 2.6.
If we prescribe the following initial conditions
$\left\{\begin{aligned} &r=x^{2}+A_{1}\phi(\xi),\\ &a=a(x,\xi)=\frac{A}{\left(1+r\right)^{3}},\quad A>0,\\ &b=b(x,\xi)=\frac{Br}{\left(1+r\right)^{6}},\quad B>0.\end{aligned}\right$. (2.23)
Then this solution blows up at $(t,\xi)=(0,\pm 1)$ in finite time $T=\tfrac{6}{A}$.
$V(t,x,\pm 1)\to\infty,\quad\tfrac{U(t,x,\pm 1)}{U(0,x,\pm 1)}\to\infty,\text{ as }t\to T=\tfrac{6}{A},x\to 0.$ (2.24)
Further more
$\left\{\begin{aligned} &\partial_{\xi}^{k}r\bigl|_{\xi=\pm 1}=0,\quad k\in\mathbb{N}\\ &\partial_{\xi}^{k}a\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}b\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}U\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}V\bigl|_{\xi=\pm 1}=0,\\ &\partial_{\xi}^{k}G\bigl|_{\xi=\pm 1}=0.\end{aligned}\right$. (2.25)
Proof of 2.6.
Since each $\xi$-derivative will introduce a factor of $\phi$ or turn $\partial_{\xi}^{k}\phi$ into $\partial_{\xi}^{k+1}\phi$, the claims in (2.25) follow from (2.22). ∎
2.6. Ridge blow-up and the full-wedge extension problem
The explicit formulas (2.20) immediately produce a blowing-up ridge background. What remains open is the extension of that ridge profile to a full classical solution of (2.11) on the wedge $x>0$, $|\xi|\leq 1$ while preserving enough ridge flatness to keep the reduced dynamics exact.
Theorem 2.7 (Explicit ridge background and localized finite-time blow-up).
Let
$a_{\mathrm{ridge}}(x):=\frac{A}{(1+x^{2})^{3}},\qquad b_{\mathrm{ridge}}(x):=\frac{B\,x^{2}}{(1+x^{2})^{6}},\qquad A,B>0,$
and let $(V_{\mathrm{ridge}},U_{\mathrm{ridge}})$ be defined by (2.20) with $(a,b)=(a_{\mathrm{ridge}}(x),b_{\mathrm{ridge}}(x))$. Then:
  1. (1)
    for every fixed $x\in\mathbb{R}$, the pair $(V_{\mathrm{ridge}}(t,x),U_{\mathrm{ridge}}(t,x))$ is smooth on $[0,T)$, where
    $T=\frac{6}{A};$
  2. (2)
    for every fixed $x\neq 0$, both $V_{\mathrm{ridge}}(t,x)$ and $U_{\mathrm{ridge}}(t,x)$ stay bounded as $t\uparrow T$;
  3. (3)
    the singularity is localized at the ridge apex $(x,\xi)=(0,\pm 1)$, in the sense that
    $V_{\mathrm{ridge}}(t,0)=\frac{6}{T-t}\to+\infty\qquad\text{as }t\uparrow T,$
    while
    $\frac{U_{\mathrm{ridge}}(t,x)}{U_{\mathrm{ridge}}(0,x)}=\frac{72}{2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}}$
    blows up like $(T-t)^{-2}$ at $x=0$.
In particular, the reduced ridge system (2.16) admits an explicit finite-time blow-up profile concentrated at the ridge apex.
Proof.
Substitute $(a,b)=(a_{\mathrm{ridge}}(x),b_{\mathrm{ridge}}(x))$ into (2.20). Since
$a_{\mathrm{ridge}}(0)=A,\qquad b_{\mathrm{ridge}}(0)=0,$
at $x=0$ we obtain
$V_{\mathrm{ridge}}(t,0)=-\frac{6\cdot 2A(tA-6)}{2(tA-6)^{2}}=\frac{6A}{6-tA}=\frac{6}{T-t},$
which blows up at $t=T=6/A$. Likewise,
$\frac{U_{\mathrm{ridge}}(t,x)}{U_{\mathrm{ridge}}(0,x)}=\frac{72}{2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}},$
and at $x=0$ this equals
$\frac{36}{(tA-6)^{2}}=\frac{T^{2}}{(T-t)^{2}},$
so the normalized $U$-profile also blows up at time $T$.
Now fix $x\neq 0$. Then $a_{\mathrm{ridge}}(x)<A$, hence
$Ta_{\mathrm{ridge}}(x)-6=\frac{6a_{\mathrm{ridge}}(x)}{A}-6<0.$
Therefore the denominator
$2\bigl(ta_{\mathrm{ridge}}(x)-6\bigr)^{2}+5t^{2}b_{\mathrm{ridge}}(x)^{2}$
stays strictly positive up to $t=T$, and both $V_{\mathrm{ridge}}(t,x)$ and $U_{\mathrm{ridge}}(t,x)$ remain bounded there. This proves the claimed localization of the singularity. ∎
The preceding theorem gives an explicit blowing-up ridge core, and in particular an explicit apex blow-up mechanism at $(x,\xi)=(0,\pm 1)$. The full background problem is to extend that ridge core to a classical wedge solution of (2.11). The next theorem isolates the apex flatness mechanism at $x=0$, while the full extension away from the apex remains a separate issue.
Theorem 2.8 (Preservation of ridge flatness at $x=0$).
Let $(U,V,G,P)$ be a classical solution of (2.11) on $0,t_{*}]\times\Omega$,
$\Omega:=\{(x,\xi):x>0,\ |\xi|\leq 1\},$
and fix a ridge $\xi_{0}\in\{\pm 1\}$. Assume that along $\xi=\xi_{0}$ one has
$G=V,\qquad U_{\xi}=V_{\xi}=G_{\xi}=P_{\xi}=0.$ (2.26)
Then the ridge flatness at $x=0$ is preserved by the PDE system (2.11).
$\left\{\begin{aligned} &\lim_{x\to 0}\bigl((\partial_{\xi}U)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\\ &\lim_{x\to 0}\bigl((\partial_{\xi}V)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\\ &\lim_{x\to 0}\bigl((\partial_{\xi}G)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0,\end{aligned}\right$. (2.27)
and
$\left\{\begin{aligned} &\lim_{x\to 0}\bigl((\partial_{\xi}P)_{t}(t,x,\xi)\bigr|_{\xi=\pm 1}\bigr)=0.\end{aligned}\right$. (2.28)
Proof of Preservation of ridge flatness at $x=0$ 2.8.
Differentiation of the $U$-equation in (2.11) with respect to $\xi$ leads to.
$\left\{\begin{aligned} \lambda(U_{\xi})_{t}&=-xU_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{1}(F,xF_{x})\cdot J_{1}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.29)
Differentiation of the $V$-equation in (2.11) with respect to $\xi$ leads to.
$\left\{\begin{aligned} \lambda(V_{\xi})_{t}&=-xV_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{2}(F,xF_{x})\cdot J_{2}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.30)
Differentiation of the $G$-equation in (2.11) with respect to $\xi$ leads to.
$\left\{\begin{aligned} \lambda(G_{\xi})_{t}&=-xG_{x}\tfrac{2\xi}{\left(\xi^{2}+1\right)^{2}}(G+V)\\ &+H_{3}(F,xF_{x})\cdot J_{3}(F_{\xi\xi},F_{x\xi},F_{\xi})\quad F\in\{U,V,G,P\}.\end{aligned}\right$. (2.31)
Here $H_{1},H_{2},H_{3}$ denote three vector functions whose elements are linear and homogeneous function of its arguments. Similarly $J_{1},J_{2},J_{3}$ denote three vector functions whose elements are linear and homogeneous function of its arguments. Since the arguments of the latter three are $\xi$-derivatives of extremely flat ridge functions, they vanish on the ridges (cf.(2.25)).
The claims in (2.27) follow after set $\xi=\pm 1$ in (2.29),(2.30),(2.31), and then take the limit of $x\to 0$.
From (2.12), we obtain
$\left\{\begin{aligned} &P_{\xi\xi}+\tfrac{2\xi}{\xi^{2}+1}P_{\xi}\\ &=-\tfrac{x\left(P_{x}+xP_{xx}\right)}{\left(\xi^{2}+1\right)^{2}}+\tfrac{2x\left(xU_{x}-\left(\xi^{3}+\xi\right)U_{\xi}\right)U(\tau,x,\xi)}{\left(\xi^{2}+1\right)^{5/2}}\\ &\quad+\tfrac{x^{2}U^{2}}{\left(\xi^{2}+1\right)^{2}}\end{aligned}\right$. (2.32)
[…]
$P_{t\xi\xi}+\tfrac{2\xi}{\xi^{2}+1}P_{t\xi}=0,$ (2.33)
Thus the solution is
$\partial_{t}(P_{\xi})(t,0,\xi)=\tfrac{C(t)}{\xi^{2}+1},$ (2.34)
Since $(P_{\xi})(0,0,\xi)=0$, we have to set $c(t)=0$ and obtain $\partial_{t}(P_{\xi})(t,0,\xi)=0$. This completes the proof of Theorem2.8.
Remark 2.9 (Interpretation of Theorem 2.8).
If Theorem 2.8 is taken as established, then the explicit ridge reduction continues to govern the apex dynamics at $(x,\xi)=(0,\pm 1)$: the first-order ridge-flatness constraints are propagated at the apex, so the pointwise ODE system (2.16) remains the correct leading-order mechanism for the blow-up there. What this theorem does not give by itself is a closed-form description of the full background for general $x>0$ and $|\xi|\leq 1$. Accordingly, the explicit part of the present theory is the apex blow-up mechanism, while the extension away from the apex remains conditional and is delegated to the background/control problem for the full wedge.
Conjecture 2.10 (Existence of a wedge background blow-up profile).
There exist a time $T>0$ and smooth functions
$U,V,G,P\in C^{\infty}\bigl([0,T)\times\Omega\bigr)$
with smooth initial data chosen compatibly with the ridge/apex symmetry class of the paper, such that the following hold.
(1) Wedge background evolution. The quadruple $(U,V,G,P)$ solves the background system (2.11) on $0,T)\times\Omega$ with the required compatibility, regularity, and first-order ridge-flatness conditions at the apex points $(x,\xi)=(0,\pm 1)$.
(2) Preservation of the apex blow-up dynamics. Along the apex trajectories $(x,\xi)=(0,\pm 1)$, the solution reduces to the distinguished ridge ODE dynamics identified in Theorem 2.7, and the corresponding finite-time blow-up law is preserved up to time $T$. In particular,
$|V(t,0,\pm 1)|\sim\frac{c_{*}}{T-t}\qquad\text{as }t\uparrow T,$
for some constant $c_{*}>0$.
(3) Critical-size background bounds. There exist constants $C_{1},C_{2}>0$ such that for all $t\in[0,T)$,
$\|V(t)\|_{L^{\infty}(\Omega)}\leq\frac{C_{1}}{T-t},\qquad\|U(t)\|_{L^{\infty}(\Omega)}\leq\frac{C_{2}}{T-t}.$
Moreover, the associated fields $G$ and the derivatives of $(U,V,G,P)$ required by the perturbative analysis obey bounds of the same critical scale.
(4) Compatibility with the perturbative bootstrap. The background coefficient bounds are strong enough to justify all weighted estimates used in the perturbation theory, so that the bootstrap assumptions can in principle be closed up to time $T$ for sufficiently small perturbations of this background.
3. Perturbation PDEs
We derive the remainder system around a prescribed background whose ridge/apex behavior is the one identified above. The full solutions are written as the sum of the background fields and the remainders:
[…]
After substituting  (3.1) into  (2.8)(1,2,3,4), we separate the full system into the background equations for $(V,U,G,P)$ and the exact remainder equations for $(v,u,g,p)$. All pure-background terms are kept in the background equations, so the remainder system contains only linear couplings to the background and genuinely nonlinear remainder–remainder interactions:
$\left\{\begin{aligned} \lambda u_{t}&=\tfrac{1}{2}(uV+Uv)-(g+v)\xi U_{\xi}-(G+V)\xi u_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xu_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xU_{x}+N_{1}\\[5.69054pt] \lambda v_{t}&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\qquad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\xi p_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x}+N_{2}+\tfrac{1}{x}p_{x}\\[5.69054pt] \lambda g_{t}&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\quad\quad\,\,\,-\tfrac{(1+\xi^{2})}{x^{2}}\tfrac{\mu}{\xi}p_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}+N_{3}+\tfrac{\mu}{x}p_{x}\\[5.69054pt] 0&=\xi xg_{x}+xv_{x}-({\xi^{2}+1})(g-v+\xi g_{\xi}+\xi v_{\xi})\end{aligned}\right$. (3.2)
Where the nonlinear perturbation terms are given by
$\left\{\begin{aligned} N_{1}:&=\tfrac{1}{2}uv\qquad\quad\,\,\,-(\xi^{2}g-v)xu_{x}-(g+v)\xi u_{\xi},\\[5.69054pt] N_{2}:&=v^{2}-u^{2}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xv_{x}-(g+v)\xi v_{\xi},\\[5.69054pt] N_{3}:&=-g^{2}\quad\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xg_{x}-(g+v)\xi g_{\xi}.\end{aligned}\right$. (3.3)
The divergence-free condition for the perturbation $(v,g)$ (3.2)(4) can be solved with the stream function $\psi(t,x,\xi)$ defined below:
[…]
3.1. Getting rid of $(p_{\xi},p_{x})$
Our next step is to get rid of $(p_{\xi},p_{x})$. Define $(\omega,\Omega)$ as
[…]
So we obtain
[…]
Now we rewrite  (3.2)2,3 as:
[…]
where $(A,B)$ are defined as
$\left\{\begin{aligned} A:&=2(vV-uU)-(g+v)\xi V_{\xi}-(G+V)\xi v_{\xi}\\ &\,\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xv_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xV_{x},\\[5.69054pt] B:&=-2gG\qquad\,\,\,-(g+v)\xi G_{\xi}-(G+V)\xi g_{\xi}\\ &\,\,-\tfrac{1}{1+\xi^{2}}(\xi^{2}G-V)xg_{x}-\tfrac{1}{1+\xi^{2}}(\xi^{2}g-v)xG_{x}.\end{aligned}\right$. (3.8)
The solutions for $(p_{x},p_{\xi})$ then become
[…]
Using $p_{x\xi}=p_{\xi x}$ to get rid of $p$, and using $g_{xt}$ of  (3.6) to simplify the result, we finally obtain:
$\left\{\begin{aligned} \tfrac{3}{2}\omega_{t}&=L_{2}+M_{2}\\[5.69054pt] L_{2}&=-\left(\tfrac{5}{3}xA_{x}+xB_{x}\right)-\left(1+\xi^{2}\right)\Bigl(\tfrac{5}{3\xi}A_{\xi}-\xi B_{\xi}\Bigr)\\[5.69054pt] M_{2}&=-\left(\tfrac{5}{3}xN_{2x}+xN_{3x}\right)-\left(1+\xi^{2}\right)\Bigl(\tfrac{5}{3\xi}N_{2\xi}-\xi N_{3\xi}\Bigr).\end{aligned}\right$. (3.10)
Substituting $(A,B)$ of  (3.8) into  (3.10)(2) and using $(g_{x},g_{xx},g_{x\xi},G_{x},G_{xx},G_{x\xi})$ of  (3.6) to simplify the results, and using (3.4) express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, we obtain
[…]
Substituting $N_{2}$, $N_{3}$ of  (3.3) into  (3.10)(3) and $(g_{x},g_{xx},g_{x\xi},G_{x},G_{xx},G_{x\xi})$ of  (3.6) to simplify the results, and using (3.4) express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, we obtain
$\left\{\begin{aligned} M_{2}&=\tfrac{10}{3\xi}\left(\left(\xi^{2}+1\right)u_{\xi}+\xi xu_{x}\right)u\\ &+\tfrac{\xi^{2}-1}{\xi^{2}+1}\left(x\psi_{x}+2\psi\right)\omega\\ &-\left(x\psi_{x}+2\psi\right)\xi\omega_{\xi}\\ &+\big(\xi\psi_{\xi}+\tfrac{1-\xi^{2}}{\xi^{2}+1}\psi\bigr)x\omega_{x}.\end{aligned}\right$. (3.12)
Using (3.4) to express $(v,g)$ in terms of $(\psi,\psi_{x},\psi_{\xi})$, the equation for $u_{t}$ in  (3.2)(1) can also be converted into desired form:
$\tfrac{3}{2}u_{t}$ $=L_{1}+M_{1},$ (3.13)
$\left\{\begin{aligned} L_{1}&=\tfrac{1}{2}V\,\,u-(G+V)\xi u_{\xi}+\tfrac{1}{\xi^{2}+1}\left(V-\xi^{2}G\right)xu_{x}\\ &+\tfrac{1}{2}\left(U-4\xi U_{\xi}-\tfrac{2}{\xi^{2}+1}\left(\xi^{2}-1\right)xU_{x}\right)\psi\\ &+\tfrac{1}{2}\left(U+2xU_{x}\right)\xi\psi_{\xi}+\tfrac{1}{2}\left(\tfrac{\xi^{2}}{\xi^{2}+1}U-2\xi U_{\xi}\right)x\psi_{x}\end{aligned}\right$. (3.14)
$\left\{\begin{aligned} M_{1}&=\tfrac{1}{2}u\psi+\tfrac{1}{2}u\left(\xi\psi_{\xi}+\tfrac{\xi^{2}}{\xi^{2}+1}x\psi_{x}\right)\\ &-\tfrac{1}{\xi^{2}+1}\psi\left(2\left(\xi^{2}+1\right)\xi u_{\xi}+\left(\xi^{2}-1\right)xu_{x}\right)\\[5.69054pt] &+\left(xu_{x}\xi\psi_{\xi}-\xi u_{\xi}x\psi_{x}\right).\end{aligned}\right$. (3.15)
Substitution of  (3.4) into  (3.5) leads to
$\left\{\begin{aligned} \omega:&=\Delta\psi,\\ \Delta:&=-\tfrac{\left(5\xi^{2}+3\right)}{3\left(\xi^{2}+1\right)}x^{2}\psi_{xx}-\tfrac{\left(19\xi^{2}+21\right)}{3\left(\xi^{2}+1\right)}x\psi_{x}-\tfrac{4}{3}x\xi\psi_{x\xi}\\ &-\tfrac{10\left(\xi^{2}+1\right)}{3\xi}\psi_{\xi}-\tfrac{1}{3}\left(\xi^{2}+1\right)\left(3\xi^{2}+5\right)\psi_{\xi\xi}\end{aligned}\right$. (3.16)
4. Explicit background, perturbation PDEs, and energy bounds
4.1. Background Solutions
The formulas below define a candidate background pair $V(t,x,\xi),U(t,x,\xi)$ associated with the end-vanishing seed profiles from (2.23). They agree with the explicit ridge reduction at the apex $(x,\xi)=(0,\pm 1)$ and are used throughout as the background ansatz for the conditional perturbative analysis:
[…]
The initial seed profiles are
$\left\{\begin{aligned} &r=x^{2}+A_{1}\phi(\xi),\\ &a(x,\xi)=\frac{A}{(1+r)^{3}},\\ &b(x,\xi)=\frac{B\,r}{(1+r)^{6}},\end{aligned}\right$. (4.2)
with $\phi$ given by (2.21). Along the ridge endpoints $\xi=\pm 1$ one has $\phi(\pm 1)=0$, hence
$r\big|_{\xi=\pm 1}=x^{2},\qquad a(x,\pm 1)=\frac{A}{(1+x^{2})^{3}},\qquad b(x,\pm 1)=\frac{B\,x^{2}}{(1+x^{2})^{6}},$
which recovers the explicit apex/ridge reduction used earlier.
Remark 4.1 (Series expansion of the new seed near the apex points).
Fix an endpoint $\xi_{0}\in\{\pm 1\}$, and write $\delta:=\xi-\xi_{0}$. Since
$\phi(\xi)=\exp\bigl(-(\xi^{2}-1)^{-2}\bigr)$
is flat at $\xi=\xi_{0}$, one has
$\phi(\xi)=O(|\delta|^{N})\qquad\text{for every }N\in\mathbb{N}$
as $\xi\to\xi_{0}$. Hence near $(x,\xi)=(0,\xi_{0})$,
$r=x^{2}+A_{1}\phi(\xi)=x^{2}+O(|\delta|^{N})\qquad\forall N\in\mathbb{N}.$
Therefore the seed profiles admit the local expansions
$a(x,\xi)=\frac{A}{(1+r)^{3}}=A-3Ar+6Ar^{2}+O(r^{3}),$
$b(x,\xi)=\frac{Br}{(1+r)^{6}}=Br-6Br^{2}+21Br^{3}+O(r^{4}),$
that is,
$a(x,\xi)=A-3Ax^{2}+O\bigl(x^{4}+\phi(\xi)\bigr),$
$b(x,\xi)=Bx^{2}+A_{1}B\phi(\xi)+O\bigl(x^{4}+x^{2}\phi(\xi)+\phi(\xi)^{2}\bigr)$
near $(0,\xi_{0})$. In particular,
$a(0,\xi_{0})=A,\qquad b(0,\xi_{0})=0,\qquad\partial_{\xi}^{k}a(0,\xi_{0})=\partial_{\xi}^{k}b(0,\xi_{0})=0\quad(k\geq 1).$
Thus the new seed is smooth and non-singular at the apex points, and its angular dependence is super-flat there. The only non-analytic feature is the standard flatness of $\phi$ at the endpoints, which is compatible with the apex reduction used in this manuscript. By contrast, the alternative seed
$r=x^{2}\bigl(1+A_{1}\phi(\xi)\bigr)$
would force $r=0$ at the origin $x=0$, so the present choice retains more transverse structure away from the apex while still giving the same leading jet at $(0,\pm 1)$.
Remark 4.2 (Seed 7 versus Seed 8 near the apex).
For the current seed choice
$\text{Seed 7:}\qquad r=x^{2}+A_{1}\phi(\xi),$
the top view near the apex displays a clear and pronounced “++” pattern centered at the apex point. For the alternative choice
$\text{Seed 8:}\qquad r=x^{2}\bigl(1+A_{1}\phi(\xi)\bigr),$
the same “++” pattern is smeared near the apex point because the angular modulation is multiplied by the vanishing factor $x^{2}$. In particular, Seed 7 preserves a sharper apex profile, while Seed 8 gives a more blended local geometry.
For the purposes of the present paper, however, both Seed 7 and Seed 8 are admissible. They produce the same apex values and the same leading-order ridge/apex jet needed for the pointwise reduction at $(x,\xi)=(0,\pm 1)$. Thus either seed can be used within the current framework, although Seed 7 gives a visually clearer apex structure.
4.2. Initial energy and finiteness
We record the “initial energy” (at $t=0$) associated with the background profile:
$E(0):=\int_{-1}^{1}\int_{-\infty}^{\infty}x^{2}\Big[a(x,\xi)^{2}+b(x,\xi)^{2}\Big]\ \frac{|x|}{1+\xi^{2}}\,dx\,d\xi.$ (4.3)
Remark 4.3 (Behavior of the weighted energy for the end-vanishing seed choice).
For the background ansatz (4.1) built from (2.23), one has
$r=x^{2}+A_{1}\phi(\xi),\qquad a(x,\xi)=\frac{A}{(1+r)^{3}},\qquad b(x,\xi)=\frac{B\,r}{(1+r)^{6}}.$
Since $0\leq\phi(\xi)\leq 1$ on [1,1][-1,1], one has $r\sim 1+x^{2}$ uniformly in $\xi$, and therefore
$a(x,\xi)=O\bigl((1+x^{2})^{-3}\bigr),\qquad b(x,\xi)=O\bigl((1+x^{2})^{-5}\bigr)\quad\text{as }|x|\to\infty,$
uniformly for $\xi\in[-1,1]$. Hence the weighted initial energy (4.3) is finite.
At the ridge endpoints one has $\phi(\pm 1)=0$, so
$a(0,\pm 1)=A,\qquad b(0,\pm 1)=0,$
and the blow-up time induced by the ridge ODE remains
$T=\frac{6}{A}.$
More generally, every $\xi$-derivative of $\phi$ vanishes at $\xi=\pm 1$, so all endpoint angular jets of $a$ and $b$ vanish there; this is the main reason the new seed family is compatible with the apex-only ridge analysis developed in the present paper.
4.3. Initial pressure determined by the PPE
At $t=0$, the explicit background satisfies
$U(0,x,\xi)=b(x,\xi)=\frac{B\,r}{(1+r)^{6}},\qquad r=x^{2}+A_{1}\phi(\xi).$
Therefore
$U(0,x,\xi)^{2}=\frac{B^{2}r^{2}}{(1+r)^{12}}=:F(r),\qquad F^{\prime}(r)=\frac{2B^{2}r(1-5r)}{(1+r)^{13}}.$
Since
$r_{x}=2x,\qquad r_{\xi}=A_{1}\phi^{\prime}(\xi),$
one obtains
$(U(0)^{2})_{x}=2x\,F^{\prime}(r),\qquad(U(0)^{2})_{\xi}=A_{1}\phi^{\prime}(\xi)F^{\prime}(r).$
Substituting these formulas into the Poisson pressure equation (2.12) yields the initial elliptic problem for
$P_{0}(x,\xi):=P(0,x,\xi):$
$P_{0,xx}+\frac{1}{x}P_{0,x}+\frac{2(1+\xi^{2})\xi}{x^{2}}P_{0,\xi}+\frac{(1+\xi^{2})^{2}}{x^{2}}P_{0,\xi\xi}$ (4.4)
$\qquad=\frac{B^{2}r^{2}}{(1+r)^{12}}-\frac{4A_{1}B^{2}}{x}\,(1+\xi^{2})^{1/2}\,\xi\,\phi^{\prime}(\xi)\,\frac{r(1-5r)}{(1+r)^{13}}+4B^{2}x(1+\xi^{2})^{-1/2}\,\frac{r(1-5r)}{(1+r)^{13}},$
where $r=x^{2}+A_{1}\phi(\xi)$. Thus the initial pressure is not prescribed independently: it is determined by the elliptic solve (4.4), together with the symmetry/decay conditions imposed on $P$ in the wedge formulation.
4.4. Updated Linear PDEs and Nonlinear Terms
The final remainder system for $(u,\omega,\psi)$ is:
[…]
4.5. Coefficient Functions
The effective elliptic operator $\Delta$ in (4.5)(3) is defined by:
$\left\{\begin{aligned} \Delta&=c_{1}(\xi)x^{2}\psi_{xx}+c_{2}(\xi)x\psi_{x}+c_{4}(\xi)x\psi_{x\xi}+c_{3}(\xi)\psi_{\xi}+c_{5}(\xi)\psi_{\xi\xi}\\ c_{1}&=-\tfrac{\left(5\xi^{2}+3\right)}{3\left(\xi^{2}+1\right)},\quad c_{2}=-\tfrac{\left(19\xi^{2}+21\right)}{3\left(\xi^{2}+1\right)}\quad c_{4}=-\tfrac{4}{3}\xi\\ c_{3}&=-\tfrac{10}{3\xi}\left(\xi^{2}+1\right)\quad c_{5}=-\tfrac{1}{3}\left(\xi^{2}+1\right)\left(3\xi^{2}+5\right).\end{aligned}\right$. (4.6)
Remark 4.4.
The apparent singular factor $\tfrac{1}{\xi}$ in $c_{3}(\xi)$ is absorbed by the weighted derivative $D_{\xi}=\tfrac{1}{\xi}\partial_{\xi}$: writing
$c_{3}(\xi)=\widetilde{c}_{3}(\xi)\,\tfrac{1}{\xi},\qquad\widetilde{c}_{3}(\theta):=-\tfrac{10}{3}\left(\xi^{2}+1\right),$
we have $c_{3}(\xi)\,\partial_{\xi}=\widetilde{c}_{3}(\xi)\,D_{\xi}$ with bounded smooth $\widetilde{c}_{3}$ on [1,1][-1,1]. Likewise, terms involving $\tfrac{1}{\xi}$ are treated as bounded multipliers in the weighted energy once expressed in terms of $D_{\xi}$ (or placed into divergence form in $\xi$).
The coefficients $c_{1}(\xi),c_{2}(\xi),c_{4}(\xi),c_{5}(\xi)$ are bounded on [1,1][-1,1]. The coefficient $c_{3}(\xi)$ contains the factor $\tfrac{1}{\xi}$ and is therefore singular at $\xi=0$ in its raw form; however, as explained above, the combination $c_{3}(\xi)\partial_{\xi}$ is naturally rewritten as a bounded multiplier times the adapted derivative $D_{\xi}=\tfrac{1}{\xi}\partial_{\xi}$. This is the form used throughout the energy estimates.
4.6. Initial conditions and boundary conditions
Boundary conditions (for the $\xi$-edges and for $x\to\pm\infty$) and initial conditions are:
[…]
The phrase “sufficiently fast decay” as $\xi\to\pm 1$ means enough vanishing and regularity near the edge so that the weighted Sobolev norms used below are finite and the boundary terms produced by integration by parts vanish at $\xi=\pm 1$.
Remark 4.5.
If we define
$\left\{\begin{aligned} &y=\log x,\quad\Delta_{1}=\Delta\\ &\omega_{1}(t,y,\xi)=\omega(t,x,\xi),\\ &\psi_{1}(t,y,\xi)=\psi(t,x,\xi),\end{aligned}\right$. (4.8)
Then $\omega=\Delta\psi$ in (4.5) becomes
[…]
Thus (4.9) becomes a well-defined elliptic problem in the strip $\Omega=\{(y,\xi):y\in\mathbb{R},\ \xi\in[-1,1]\}$.
5. Linear estimates and conditional nonlinear control up to blow-up time
Updated perturbation system. Throughout this section we work with the final perturbation equations derived in Section 2. We study the perturbation system (3.13) and (3.10), together with the coefficient collections (3.14),(3.15),(3.11),(3.12) and the elliptic operator (4.6), on the time interval $[0,T)$ up to the background blow-up time, around the explicit background (4.1). Throughout, all Lebesgue and Sobolev norms are taken with respect to the weighted measure $d\mu_{w}=w(\xi)|x|\,dx\,d\xi$, and we use the desingularized angular derivative
$D_{\xi}:=\tfrac{1}{\xi}\partial_{\xi}.$
This is the natural derivative after the change of variables $\xi=\tan\theta$, which turns the trigonometric coefficients into rational functions of $\xi$.
5.1. Bootstrap framework and adapted background coefficient bounds
Fix an integer $k\geq 6$. Define the perturbation energy
$\mathcal{E}_{k}(t):=\sum_{\begin{subarray}{c}j+\ell\leq k\end{subarray}}\Big(\|\partial_{x}^{j}D_{\xi}^{\ell}u(t)\|_{L^{2}_{\mu_{w}}}^{2}+\|\partial_{x}^{j}D_{\xi}^{\ell}\omega(t)\|_{L^{2}_{\mu_{w}}}^{2}\Big)+\sum_{\begin{subarray}{c}j+\ell\leq k+1\end{subarray}}\|\partial_{x}^{j}D_{\xi}^{\ell}\psi(t)\|_{L^{2}_{\mu_{w}}}^{2}.$ (5.1)
Background coefficient bounds actually needed in the energy method. The candidate background (4.1) provides a closed-form model for $(U,V)$ and, in particular, reproduces the explicit apex dynamics. What the stability estimates require is not a uniform bound on the raw derivatives $\partial_{x}^{m}\partial_{\theta}^{\ell}V$ (which can grow faster than $(T-t)^{-1}$ near the intermediate scale $r^{2}\sim T-t$), but rather uniform control of the degenerate combinations that appear in (3.14),(3.11),(3.12) and in the weighted Sobolev norms.
Define the adapted derivatives
$Z_{x}:=x\partial_{x},\qquad D_{\xi}:=\tfrac{1}{\xi}\partial_{\xi},$
Then for each integer $k\geq 0$ there exists $C_{*}=C_{*}(A,B,\text{seeds},k)$ such that for all $t\in[0,T)$ the following estimate holds.
Lemma 5.1 (Adapted background coefficient bounds).
$\sum_{j+\ell\leq k}\Big(\|Z_{x}^{j}D_{\xi}^{\ell}V(t)\|_{L^{\infty}}+\|Z_{x}^{j}D_{\xi}^{\ell}U(t)\|_{L^{\infty}}\Big)$ (5.2)
$+\sum_{j+\ell\leq k}\Big(\Big\|Z_{x}^{j}D_{\xi}^{\ell}\bigl(x\,V_{x}(t)\bigr)\Big\|_{L^{\infty}}+\Big\|Z_{x}^{j}D_{\xi}^{\ell}\bigl(x\,U_{x}(t)\bigr)\Big\|_{L^{\infty}}\Big)$
$\;\leq\;\frac{C_{*}}{T-t}.$
Proof sketch.
The estimate is a direct consequence of the closed-form background formula together with the seed expansions near the ridge apex and the fact that away from the apex the denominator in (4.1) is bounded below uniformly. Near the apex one uses the denominator scale $D(t,x,\xi)\gtrsim(T-t)^{2}+r^{2}$, together with the observation that $Z_{x}=x\partial_{x}$ acts like $r\partial_{r}$ and $D_{\xi}$ neutralizes the harmless trigonometric singular factors in angular derivatives. ∎
In particular,
$\|V(t)\|_{L^{\infty}}=\frac{6}{T-t}\quad\text{(attained at }(x,\xi)=(0,\pm 1)\text{)},\qquad\|U(t)\|_{L^{\infty}}\lesssim\frac{1}{T-t}.$
5.2. Elliptic control of $\psi$ from $\omega=\Delta\psi$
The elliptic relation in (4.5) reads $\omega=\Delta\psi$, where $\Delta$ is given by (4.6). After rewriting the angular part in terms of the adapted derivative $D_{\xi}$ (as already indicated in the weighted-norms subsection), the operator $\Delta$ has the same principal structure as $\Delta_{1}=\partial_{\log x}^{2}+D_{\xi}^{2}$ with lower-order $\xi$-dependent coefficients controlled on the wedge. Accordingly, we record the following weighted elliptic estimate as the analytic input needed for the perturbation argument: for all integers $m\geq 0$,
$\|\psi(t)\|_{H^{m+2}_{\mu_{w}}}\leq C_{\Delta,m}\,\|\omega(t)\|_{H^{m}_{\mu_{w}}},$ (5.3)
where the constant depends only on the wedge geometry, the boundary conditions, and the bounded coefficient functions appearing in (4.6). This estimate is natural from the $y=\log x$ reformulation discussed in Remark 4.5; in the present manuscript we use it as a working elliptic input for the $\psi$-estimate rather than as a separately proved theorem.
In particular, since $k\geq 6$, Sobolev embedding in the $(x,\xi)$ variables (with $D_{\xi}$ counted as one derivative) gives
$\|u(t)\|_{L^{\infty}}+\|\omega(t)\|_{L^{\infty}}+\|\psi(t)\|_{W^{1,\infty}}\leq C\,\mathcal{E}_{k}(t)^{1/2}.$ (5.4)
5.3. Energy inequality for $(u,\omega)$
Differentiate the $u$-equation and the $\omega$-equation in (4.5) by $\partial_{x}^{j}D_{\xi}^{\ell}$ for $j+\ell\leq k$, take the $L^{2}_{\mu_{w}}$ inner product with $\partial_{x}^{j}D_{\xi}^{\ell}u$ and $\partial_{x}^{j}D_{\xi}^{\ell}\omega$, and sum over $j+\ell\leq k$. The transport terms are now written directly in the $(x,\xi)$ variables, so the integration-by-parts step is carried out in $x$ and $\xi$. The boundary contributions vanish because of the remainder boundary conditions at $\xi=\pm 1$, the decay as $|x|\to\infty$, and the weighted formulation using $D_{\xi}=\xi^{-1}\partial_{\xi}$.
Using the commutator estimates and (5.4), one obtains an inequality of the form
$\frac{d}{dt}\mathcal{E}_{k}(t)\leq\frac{C_{\rm lin}}{T-t}\,\mathcal{E}_{k}(t)+C_{\rm nl}\,\Big(\|M_{1}(t)\|_{H^{k}_{\mu_{w}}}+\|M_{2}(t)\|_{H^{k}_{\mu_{w}}}\Big)\,\mathcal{E}_{k}(t)^{1/2}.$ (5.5)
Quadratic remainder terms. From the explicit forms of $M_{1},M_{2}$ in (3.15) and (3.12), together with Moser and Sobolev product estimates in the $(x,\xi)$ variables, one obtains
$\|M_{1}(t)\|_{H^{k}_{\mu_{w}}}+\|M_{2}(t)\|_{H^{k}_{\mu_{w}}}\leq C\,\mathcal{E}_{k}(t).$
Because all pure-background terms have been kept in the background system, there is no additive forcing term in the remainder energy inequality. Thus it is natural to rewrite (5.5) in terms of
$Y(t):=\mathcal{E}_{k}(t)^{1/2}.$
Then
$Y^{\prime}(t)\leq\frac{C_{\rm lin}}{T-t}Y(t)+C_{\rm nl}Y(t)^{2}$ (5.6)
whenever $Y(t)>0$.
The important point is that (5.6) by itself does not yet imply a closed bootstrap with a remainder strictly smaller than the background singularity. What it does give is an Elgindi-type conditional transfer principle: if the remainder stays in a class whose growth is weaker than the background blow-up rate, then the quadratic term is perturbative and the background singularity transfers to the full solution.
To make this precise, fix an exponent $\sigma>0$ and define the renormalized energy envelope
$X_{\sigma}(t):=(T-t)^{\sigma}Y(t).$
Differentiating and using (5.6) gives
$X_{\sigma}^{\prime}(t)\leq\frac{C_{\rm lin}-\sigma}{T-t}X_{\sigma}(t)+C_{\rm nl}(T-t)^{-\sigma}X_{\sigma}(t)^{2}.$ (5.7)
Hence, whenever
$\sigma>C_{\rm lin},$ (5.8)
and whenever a bootstrap bound of the form
$X_{\sigma}(t)\leq M\varepsilon\qquad\text{for }0\leq t\leq t_{*}$ (5.9)
holds with $\varepsilon>0$ sufficiently small, the right-hand side of (5.7) is integrable and the quadratic term can be absorbed. Standard continuity then yields
$X_{\sigma}(t)\leq 2X_{\sigma}(0)\qquad\text{for }0\leq t\leq t_{*}.$ (5.10)
Equivalently,
$Y(t)\leq 2Y(0)\Bigl(\frac{T}{T-t}\Bigr)^{\sigma},\qquad 0\leq t\leq t_{*}.$ (5.11)
In particular, if one can choose $\sigma<1$ while still having (5.8), then the remainder stays strictly below the background blow-up scale $(T-t)^{-1}$ in the detecting norm.
This discussion is summarized in the following conditional theorem.
Theorem 5.2 (Conditional nonlinear control up to the background blow-up time).
Assume that a compatible background solution exists on $[0,T)$, has the same apex blow-up rate as the explicit ridge dynamics at $(x,\xi)=(0,\pm 1)$ with time $T=6/A$, and satisfies the adapted coefficient bounds of Lemma 5.1. Assume also that the weighted elliptic estimate (5.3) holds. Let $k\geq 6$, and let $(u,\omega,\psi)$ solve the exact remainder system on $0,t_{*}]\subset[0,T)$.
Then there exist constants $C_{\rm lin},C_{\rm nl}>0$, depending only on $k$ and the background coefficient bounds, such that (5.6) holds. Consequently, for every exponent $\sigma$ satisfying (5.8), there exists $\varepsilon_{0}=\varepsilon_{0}(\sigma,k)>0$ with the following property: if
$X_{\sigma}(0)=T^{\sigma}\mathcal{E}_{k}(0)^{1/2}\leq\varepsilon_{0},$
and if the bootstrap assumption (5.9) holds on $0,t_{*}]$, then in fact (5.10) holds on $0,t_{*}]$.
Remark 5.3 (What this proves now, and what still has to be improved).
Theorem 5.2 is already strong enough to put the remainder analysis into the same logical class as the Elgindi–Jeong mechanism: the singular core is the explicit background, and the nonlinear argument reduces to showing that the remainder remains in a better class. However, the theorem is still conditional. To turn it into a full stability statement one still needs an independent argument guaranteeing a gap (5.8) with some exponent $\sigma<1$ in the norm that detects the background blow-up. This may come from sharper coercivity, additional vanishing of the remainders at the ridge, or a more scale-adapted energy functional.
Under this conditional control, one obtains a blow-up transfer statement for the full solution.
Theorem 5.4 (Conditional transfer of background blow-up to the full solution).
Assume the hypotheses of Theorem 5.2. In addition, suppose that the chosen detecting norm $\mathcal{N}_{\rm det}(t)$ for the full solution satisfies
$\mathcal{N}_{\rm det}^{\rm bg}(t)\sim c_{0}(T-t)^{-1}\qquad\text{for some }c_{0}>0,$
when evaluated on the background, and that the remainder contribution is estimated by
$\mathcal{N}_{\rm det}^{\rm rem}(t)\leq C_{\rm det}Y(t)$
for $0\leq t<T$. If there exists $\sigma\in(C_{\rm lin},1)$ such that (5.10) holds on $[0,T)$, then
$\mathcal{N}_{\rm det}(t)=\mathcal{N}_{\rm det}^{\rm bg}(t)+O\bigl((T-t)^{-\sigma}\bigr)\qquad\text{as }t\uparrow T,$
and hence the full solution blows up at time $T$ with the same leading-order singularity location and blow-up scale as the background.
Accordingly, the logical bottleneck of the manuscript is no longer a forcing obstruction in the remainder equations. The main unresolved issue is instead the rigorous construction/control of a background away from the apex, with the coefficient bounds needed by the weighted energy method and with enough rigidity near the apex to match the explicit ridge dynamics, together with whatever refined estimate is needed to produce a genuine gap exponent $\sigma<1$ in the remainder norm. Once those two inputs are available, the present stability mechanism upgrades directly to a nonlinear remainder theorem in the spirit of Elgindi.
6. Conclusion
We derived a closed (1+2)(1+2)D subsystem (E1) from the inviscid 2D Boussinesq equations under a parity ansatz and organized its blow-up analysis around two components: an explicit ridge/apex core and an exact remainder system. In this formulation, the ridge rays carry a convection-free (1+1)(1+1)D CLM-type reaction dynamics with explicit finite-time blow-up at the apex $(x,\xi)=(0,\pm 1)$, while the full wedge problem is rewritten in the $(x,\xi)$ variables so that the remainder equations are exact and all pure-background terms remain in the background system.
The weighted energy method developed in Section 5 shows that, if a compatible background exists on $[0,T)$ with the coefficient bounds required there and with apex trace governed by the explicit ridge dynamics, and if the remainder stays subordinate to the background singularity in the detecting norm, then the full solution inherits the same finite-time blow-up.
The main unresolved step is therefore the construction and control of a full background away from the apex, together with the rigidity properties needed to match the apex dynamics and close the nonlinear bootstrap without loss. The blow-up mechanism itself is available in closed form at the apex, but extending that information to a full background with the necessary compatibility bounds remains the decisive open problem.
Even before the final nonlinear theorem is completed, the present formulation already isolates the core components of the analysis. It provides an exact derivation from Boussinesq, an explicit apex blow-up mechanism, a strong linearized stability framework, conditional nonlinear control, and a precise conditional blow-up transfer statement.
Natural next steps are therefore clear. The first is to prove the full background existence/control theorem compatible with the apex dynamics identified here. The second is to sharpen the detecting norm so that the remainder remains strictly below the background blow-up rate, yielding a closed nonlinear bootstrap. After that, one can revisit modulation of geometric parameters and lower-regularity weighted theories.
7. Acknowledgements
ChatGPT is credited as a substantive contributor to drafting and technical editing; responsibility for correctness remains with the author. The author thanks Prof. Zixiang Zhou of the Department of Mathematics at Fudan University and Prof. Jie Qin of the Department of Mathematics at the University of California, Santa Cruz for their continuous support and encouragement over the years. The author also thanks colleagues and the broader PDE and fluid-dynamics community for stimulating discussions on inviscid Boussinesq and CLM-type models. (Computational assistance: ChatGPT, GPT–5.2 Thinking,GPT–5.4 Thinking, OpenAI; sessions in March 2026.)